### v1

# CrowdGenUI: Enhancing LLM-Based UI Widget Generation with a Crowdsourced Preference Library

Abstract.
Large Language Models (LLMs) have demonstrated remarkable skills across various design domains, including UI generation. However, current LLMs for UI generation tend to offer generic solutions that lack a deep understanding of task context and user preferences in specific scenarios. We present CrowdGenUI, a framework that enhances LLM-driven UI generation with a crowdsourced user preference library. This approach addresses the limitations of existing methods by guiding LLM reasoning with user preferences, enabling the generation of UI widgets that align more closely with user needs and task-specific requirements. Using image editing as a test domain, we built this library from 50 users, capturing 720 user preferences, which include the predictability, efficiency, and explorability of multiple UI widgets. In a user study with 72 additional participants, our framework outperformed standard LLM-generated widgets in meeting user preferences and task requirements. We discuss these findings to inform future opportunities for designing user-centered and customizable UIs by comprehensively analyzing the extendability of the proposed framework and crowdsourced library.
user preference, UI generation, large language models
$c$s: Human-centered computing Collaborative and social computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing User studiesccs: Information systems Crowdsourcing
Figure 1. CrowdGenUI is a framework that allows the generation of task requirement-aware and user preference-aligned UI widgets powered by LLMs. Based on input information specifying the content to edit, the task to perform, and various user preferences, the LLM-driven UI widget generation is enhanced by our crowdsourced user preference library together with a set of widget candidates for UI widget reasoning and code generation.
1. Introduction
Graphical user interfaces (GUIs) play a fundamental role in how most people operate computers and other computing devices. Carefully designed by expert interface designers, these GUIs aim to provide users with learnable ways to interact with software (Adobe, 2024c, b). Although effective, such interfaces can be complicated and sometimes overwhelming for users. In recent years, the emergence of large language models (LLMs) has led to increased interest in natural language interfaces. These language-based interfaces process human language input, analyze high-level intentions, and turn the interpreted intents into operations, providing a simplified way for users to interact with technology (OpenAI, 2024a; Descript, 2024; Maddigan and Susnjak, 2023). However, while language interfaces are promising, they are typically suited for high-level task execution, such as removing an object from an image or adding a filtering effect to videos. They often lack the precision required when users need to fine-tune specific parameters, such as adjusting the color in an image or customizing detailed settings in a video editing project.
To address the complexity of professional GUIs and the inaccuracy of natural language interfaces, researchers have begun to explore the integration of LLMs with UI generation to combine the high-level task understanding enabled by LLMs with the precise, granular control offered by GUIs (Vaithilingam et al., 2024; Cheng et al., 2024). By turning natural language inputs into interactable UI widgets through UI generation, these systems provide an exciting hybrid interaction model that supports both user interaction with widgets for precise control and language as input for high-level task interpretation. Nevertheless, the existing UI generation systems also come with challenges. These LLM-generated UIs tend to be generic and may not fully account for the nuances of domain-specific tasks or the diverse preferences of individual users. Without specialized knowledge of interface design or a deeper understanding of user preferences, these UIs may fail to provide the optimal interaction experience.
To overcome these challenges, we propose a novel framework, CrowdGenUI, that enhances LLM-driven UI generation by incorporating a crowdsourced user preference library, as illustrated in Figure 1. This framework leverages the collective insights of users to guide the LLM in generating tailored UI widgets for different user preferences. While the framework is broadly applicable across various domains, we demonstrate its effectiveness in the context of image editing within a Python programming environment. To build the preference library, we collected data from 50 users, with no requirement on UI design or programming experience, capturing their preferences across three key dimensions of interface design, i.e., predictability, efficiency, and explorability. By integrating this library into the LLM’s UI widget generation process, our system generates widgets that are not only more precise but also better aligned with user preferences and task-specific requirements.
In our user study with a different set of 72 participants, we found that UI widgets generated based on our proposed framework were more aligned with user preferences and task requirements than those generic widgets generated by the LLM alone. This demonstrates the potential of our framework to improve LLM-based UI generation and suggests that crowdsourced user preference data can play a crucial role in creating user-centered user interfaces.
To summarize, this paper makes the following key contributions:
  • We propose a general framework to enhance LLM-driven UI generation by incorporating a crowdsourced user preference library for task-aware and preference-aligned widget generation.
  • As a concrete implementation of this framework, we develop a system for the domain of image editing within a Python programming environment. This system includes crowdsourced data from 50 users, capturing user preferences for UI widget configurations in image editing tasks.
  • We conduct a user study with an additional 72 participants to evaluate the effectiveness of the implemented system. The results of this study are extensively discussed to inform future research on the development of customizable and user-centered UIs.
2. Related Work
To provide the background for our work, we introduce prior work on traditional fixed UI, natural language UI, and dynamically generated UI. We discuss their advantages and limitations and compare them with our proposed framework. A visual comparison is presented in Figure 2.
2.1. Traditional Fixed UI
Traditional UI design in professional digital software like Illustrator (Adobe, 2024b), Photoshop (Adobe, 2024c), Premiere Pro (Adobe, 2024e), and Blender (Blender, 2024) follows a designer-centric model, where a fixed set of tools and controls are designed for various tasks. These applications rely on well-established interface elements such as dropdown menus, sliders, text fields, and panels that allow users to manipulate images, graphics, videos, and 3D models. For example, Photoshop provides consistent tools for adjustments like brightness, contrast, and hue, while Illustrator offers vector editing through pre-defined tools. Premiere Pro delivers video editing features through timelines and control panels, and Blender offers a complex interface for 3D modeling and animation. These tools are powerful but often rigid, requiring users to adapt to the software’s structure. As a result, their complexity can be overwhelming to users, and their interface may not always align with individual workflows or preferences, which limits the flexibility and personalization of user experience.
$2.$ Natural Language UI
Using natural language as input to edit digital content has gained increasing interest due to its simplicity. This approach allows users to perform tasks like image, video, data visualization, 3D models, and human motion editing by simply describing the changes they want rather than navigating complex software interfaces. For example, natural language-based image editing tools, such as DALL-E (OpenAI, 2024a) or Runway (Runaway, 2024), allow users to generate and edit images by typing commands like “make the sky bluer” or “add a tree in the background”. Similarly, tools like Adobe Firefly (Adobe, 2024a) and Photoshop Generative Fill (Adobe, 2024d) enable image generation and edits using prompts like “change to a vintage look” or “add a yellow bee in the circled region next to the flower”. In video editing, tools like Descript (Descript, 2024) allow users to edit video content by editing a transcript, and tools like Runway and Premiere Pro with Sensei AI (Adobe, 2024f) also support video editing using natural language prompts like “add a motion blur effect”. Also, there are many research projects (Brooks et al., 2023; Zhang et al., 2024; Fu et al., 2024, 2022; Qin et al., 2024) that have significantly advanced the techniques for image and video editing using language as the primary modality. In the field of visualization, systems like Chat2VIS (Maddigan and Susnjak, 2023), NL4DV (Sah et al., 2024) and Data Formulator (Wang et al., 2023, 2024) allow users to create charts and graphs by simply asking for a “bar chart showing sales over time”, and LIDA (Dibia, 2023) enable the generation of grammar-agnostic visualizations and infographics using LLMs. For the generation and editing of 3D model (Chang et al., 2017; Poole et al., 2022) and human motion (Tevet et al., 2022; Raab et al., 2023; Chen et al., 2023; Liu and Sra, 2024), prior research has investigated using language to describe users’ goals, such as “create a 3D lemur taking notes in a journal” or “change the upper body motion to throwing a ball and keep the original running motion”. These innovations are transforming traditional workflows by making advanced editing more accessible to non-expert users, reducing the technical execution gulf and enabling them to focus on creative intent.
However, natural language interfaces struggle with precise control over specific details and cannot effectively support interactive editing, limiting their ability to handle complex, fine-grained adjustments typically required in professional workflows.
2.3. Dynamically Generated UI
Recent research has studied generating dynamic UIs to provide users with precise, direct interactions for content editing. One prominent approach is the relaxation method, which involves creating UI widgets that provide generalized control over variables within a function or query. For instance, Mavo (Verou et al., 2016) provides users with the capability to construct interactive HTML pages by incorporating unique attributes, which, in turn, recommend editing widgets tailored to those attributes. Heer et al. (Heer et al., 2008) employ query relaxation techniques, allowing users to broaden their selection effectively. Bespoke (Vaithilingam and Guo, 2019) utilizes user demonstrations to generate custom UIs for command-line applications. It applies rule-based heuristics to infer the semantic types of parameters in bash commands and thus dynamically create widgets for parameter manipulation. Another line of research, known as precision interfaces, uses SQL queries as a framework for UI generation (Zhang et al., 2018; Chen, 2020; Chen and Wu, 2022). These systems convert a sequence of input queries into interactive widgets, allowing users to adjust query parameters directly. An example is NL2Interface (Chen et al., 2022), which translates natural language commands into SQL queries and subsequently generates a UI for users to edit parameters within the queries.
In addition to generating dynamic UIs from command-line applications and queries, recent advancements have focused on creating UIs directly from natural language descriptions. This approach allows users to articulate their goals in natural language, and the system generates an interactive UI tailored to the task at hand using LLM’s reasoning capabilities. For instance, DynaVis (Vaithilingam et al., 2024) generates dynamic interfaces for visualization editing based on natural language commands. By enabling users to describe changes to their visualizations through simple text, such as “rotate the x-labels by 30 degrees counter-clockwise”, DynaVis dynamically creates suitable widgets, such as a slider for adjusting the rotation of x-labels. Similarly, Biscuit (Cheng et al., 2024) generates UIs to help users follow machine learning tutorials in Jupyter Notebooks using natural language queries. Users can describe tasks like “sample training data” or “adjust model hyperparameters”, and Biscuit generates ephemeral UIs for the specified parameter adjustments.
Although natural language-based UI generation generates UIs without depending on fixed rules or query relaxation techniques and utilizes LLMs’ advantage to interpret intents from high-level descriptions, most prior work using standalone LLMs does not fully consider individual user preferences and domain-specific task requirements. This can result in interfaces that, while functional, may not be the most effective or usable for specific users, especially when personalization and task-specific requirements are needed.
3. User Feedback Collection and Principles for UI Design
While LLM-powered UI generation offers many advantages, it primarily relies on generic training data, lacking task-specific contexts and user-centered insights, such as individual preferences. User preference is part of user feedback and plays a crucial role in UI design and UI widget selection. The user feedback data is traditionally collected through labor-intensive methods like interviews, usability tests, A/B tests, and feedback forms (Stone et al., 2005; Wilson, 2013). Given that LLMs are trained on large-scale human data, they provide a promising solution to incorporating user preferences into dynamically generated UIs. However, without dedicated user preference datasets, current LLMs cannot fully address this need. To fill this gap, we propose crowdsourcing a preference library to enhance LLM-driven UI widget generation with user preference data, motivated by the effectiveness of adopting crowdsourcing for user feedback collection to assist various design areas (Xu et al., 2014, 2015; Luther et al., 2015; Yu et al., 2016; Foong et al., 2017; Lee et al., 2018; Oppenlaender et al., 2020).
To determine the specific aspects of user preference, we refer to well-established UI design principles developed by prior research. McKay’s UI is Communication (McKay, 2013) emphasizes that an intuitive interface relies on effective communication that balances multiple principles, including predictability, efficiency, discoverability, and explorability. McKay argues that intuitive UIs align with users’ prior experiences, providing clear affordances and responsive feedback to guide interaction. Lidwell et al.’s Universal Principles of Design (Lidwell et al., 2010) expand on these ideas by highlighting fundamental design principles such as the flexibility-usability tradeoff, mapping, and mental models, which underscore the importance of aligning system controls with user expectations and simplifying complex interactions. Together, these principles highlight that successful UIs are predictable and efficient and allow for exploratory interaction without overwhelming the user. Drawing from these foundational principles, we have identified predictability, efficiency, and explorability as three critical aspects to guide our approach to building a user preference library to enhance LLM-powered UI widget generation.
The remainder of the paper is organized as follows: first, we provide an overview of our proposed framework. Next, we describe the implemented system in the image editing domain and discuss the crowdsourced data collection process. This is followed by a detailed user evaluation, along with a discussion of the limitations, potential future work, and the conclusion.
Figure 2. Overview of our proposed CrowdGenUI framework and comparison with prior work. Our framework consists of four main components: it takes (1) task description in natural language as input, and this information is passed to the LLM together with (2) the collected user preference library and (3) a collection of UI widget candidates. All the input data is used to guide the LLM reasoning process to obtain (4) generated UI widgets that are aligned with user preferences and task requirements.
4. Framework Overview
Figure 2 provides an overview of our proposed framework, CrowdGenUI, and a visual comparison with previous research. Our framework improves existing UI widget generation systems that take task descriptions as input for LLMs to generate UIs dynamically. To align generated UI widgets with specific user needs, we collect a user preference library that captures users’ preferred UI widgets and their rationale. This library is provided to the LLM alongside task descriptions and widget candidates to guide the UI generation process. For implementation, the widget candidates are selected from the widget list of Jupyter Widgets Library (Jupyter, 2024b). These candidates can be easily extended to include additional widget options, such as those from other widget libraries or LLM-generated widgets.
5. Crowdsourced Library-enhanced UI Widget Generation for Image Editing
Figure 3 illustrates the system we implemented for our proposed framework. In this section, we dive into the technical details of this system, including UI widget reasoning, integration of the crowdsourced library into the reasoning process, and UI widget code generation.
[…]
5.1. UI Widget Reasoning
[…]
5.2. Integration of Crowdsourced User Preference Library
[…]
5.3. UI Widget Code Generation
[…]
5.4. Implementation Details
We use GPT-4o (OpenAI, 2024b) as the LLM for both widget reasoning and code generation, with Python as the programming language to implement the widgets. Occasionally, the code generated by GPT-4o may have minor bugs and might not compile correctly. We manually fix these issues to ensure all code generated by the LLM runs properly during user evaluation. This setup aligns with our focus on LLM-driven widget reasoning and code generation rather than real-time UI generation. We use this system to serve as a prototype to demonstrate how LLM reasoning, when enhanced by a crowdsourced library, can effectively guide UI widget generation according to task requirements and user preferences, particularly in the image editing domain.
6. Crowdsourcing a Preference Library for UI Widgets
In this section, we introduce how we collect user preference data to build the library that enhances LLM-driven UI widget generation. Following our previous implementation, we gather preference data in the image editing domain to illustrate the process. We present the study setup and provide an analysis of the collected data.
(a) Task description and provided UI widgets to perform the task.
(b) UI widget preference selection panel.
Figure 4. Crowdsourcing user interface for task: adjusting an image to fall colors (image_adjust_fall_color).
6.1. Platform and Participants
[…]
6.2. Crowdsourcing Setup
6.2.1. User Interface
In the crowdsourcing study, we offered each participant a Jupyter Notebook to work on eight image editing tasks, interact with several UI widgets to perform each task and provide their preferences for the widgets. Figure 4 shows the user interface for a crowdsourcing task—adjusting an image to fall colors. The interface contains the task requirement, multiple UI widgets, an image to edit, and a user preference selection panel.
6.2.2. Tasks
The crowdsourcing tasks cover three categories in the image editing domain. We introduce these categories and provide an overview of the tasks below. Some tasks may overlap across multiple categories. Our goal is to use these tasks to incorporate a range of value adjustments and data manipulation needs in image editing.
  • Continuous value adjustment: These tasks need to adjust a value across a continuous range smoothly. Examples include adjusting image lightness, saturation, hue, and color balance.
  • Discrete value selection: These tasks involve selecting from predefined options. Examples include choosing predefined values for image lightness, saturation, hue, and watermark positions.
  • Tasks emphasizing visual feedback: These tasks require immediate visual feedback, such as color or position changes. Examples include changing an image to fall tones, matching an object’s color to a given color, positioning a watermark, and highlighting a specific image area.
The detailed task descriptions are presented in Table 1. We use these tasks to encourage exploration, such as in Tasks 1, 2, 3, 6, and 7, while also emphasizing precision, as seen in Tasks 4, 5, and 8. Two tasks are used as control tasks: image_adjust_fall_color and image_color_match. We instruct participants to ensure that the final images best meet the task requirements and use these tasks as attention checks to help identify untrustworthy responses. We randomize the order of tasks, ensuring that no particular task appears consistently in the same position to reduce order effects, such as fatigue or priming, that could influence user preferences and behaviors.
Table 1. Crowdsourcing task names and descriptions.
Task Name Task Description
1 image_adjust_lightness Experiment with lightness settings to see how different levels affect the image.
2 image_adjust_saturation Boost the saturation to make the colors pop.
3 image_adjust_hue Experiment with different hues to find a color tone that complements the overall mood of the image.
4 image_adjust_fall_color Adjust the color settings to gradually shift the image tones, creating a warm, autumnal atmosphere. Ensure the final image is the one that looks the most autumnal to you.
5 image_color_match Change the rocket’s color to match the provided reference color. Ensure the rocket’s color in the final image best matches the reference color.
6 image_adjust_color_balance Experiment with different color balance settings to see how altering the red, green, and blue levels affects the overall color harmony of the image.
7 image_place_watermark Experiment with different watermark positions to find the perfect balance between visibility and subtlety.
8 image_place_vignette Darken the background using a vignette effect except for the human face by properly positioning the circle around the face.
6.2.3. User Preference Aspects
We use three key aspects of UI widget preferences, drawing from prior research on intuitive user interface design principles (recall Section 3). These aspects—predictability, efficiency, and explorability—reflect diverse user preferences when interacting with digital software, with the detailed descriptions below:
  • Predictability: Users can achieve results without encountering unexpected outcomes or needing to figure out how to perform the interaction.
  • Efficiency: Users can perform tasks with minimal effort.
  • Explorability: Users can explore multiple possibilities and perform tasks with high flexibility.
6.2.4. UI Widget Selection and Implementation
UI widget selection needs to consider task and data manipulation requirements, such as input and output data types, and match the widget types with the specified task requirements. We choose UI widgets for the crowdsourcing tasks from the Jupyter Widgets Library (Jupyter, 2024b), which offers both basic and advanced UI widgets, covering a wide range of control needs. Specifically, for continuous value adjustment tasks, sliders [Uncaptioned image] are typically ideal to allow for fine-grained control, and text fields [Uncaptioned image] can be used for precise value input. For discrete value selection tasks, dropdown menus [Uncaptioned image] and radio buttons [Uncaptioned image][Uncaptioned image] are often effective. For tasks emphasizing visual feedback, UI widgets like color pickers [Uncaptioned image], color wheels [Uncaptioned image], direct manipulation (e.g., clicking [Uncaptioned image]), and preset buttons with preview overlays [Uncaptioned image][Uncaptioned image] allow users to see outcome previews through color or position visual cues.
In this work, we present one approach to UI widget implementation and selection, acknowledging that it does not cover all possible options. However, this method can be easily expanded by leveraging additional UI widget libraries, such as (Material, 2024; Bootstrap, 2024; Vuetify, 2024). These libraries can facilitate data manipulation and content editing in software or websites and enable the inclusion of a broader range of widget types in various applications.
6.3. Crowdsourced UI Widget Preference Library
6.3.1. Overview
We distributed the crowdsourcing study to the 50 participants in three versions. Each version included all eight tasks and a subset of the three preference aspects, covering one, two, or all three aspects. This approach ensured that each participant provided their preferences for 15-18 preference selections out of the total 3 aspects $\times$× 8 tasks = 24 selections across all tasks, helping to prevent fatigue. The study yielded a UI widget preference library with 30 unique user responses for each aspect. In total, we obtained 30 responses $\times$× 3 aspects $\times$× 8 tasks = 720 user preferences. Figures 5 and 6 show the distribution of user preferences for all crowdsourcing tasks.
Figure 5. User preference distribution of tasks 1-4 collected in the crowdsourcing study.
Figure 6. User preference distribution of tasks 5-8 collected in the crowdsourcing study.
6.3.2. Analysis of User Preference Distribution
From the crowdsourcing results of all 30 responses to each preference aspect, we observe variations in user preferences for UI widgets regarding predictability, efficiency, and explorability across the eight tasks.
For continuous value adjustment tasks, including adjusting image lightness, saturation, and hue, many users preferred preset buttons with preview overlays for their predictability and efficiency, with reasons like “The preset button’s background allows for users to figure out the lightness relatively quickly, and when clicking on the buttons, the button updates the lightness to how it looks on the buttons, allowing for users to need minimal deduction to understand what this button does”. For explorability, sliders were favored for adjusting lightness and saturation. Participants commented that “The slider allows users to slowly go through each increment and explore what effect they have on the image. They can go anywhere within the range to understand the effect of saturation on the image and precisely pinpoint which exact value of saturation they want with a slider”. A color wheel was preferred to explore hue adjustment, with comments like “The color wheel, due to it allowing you to select exactly what you want and visualizing it, allowing you to explore and experiment very easily”.
For discrete value selection tasks, including choosing predefined values for image lightness, saturation, hue, and watermark positions, preset buttons were favored over radio buttons and dropdown menus in all three preference aspects. This was primarily due to the visual cues that preset buttons provided, such as user-provided reason saying that “The preset buttons include a preview of how the saturation would look like. For example, -1.0 is fully gray, while 1.0 is a bright rainbow. This allows users who may not know what saturation means to understand what each button does; it changes how colorful the image is”.
For tasks emphasizing visual feedback, preferences varied based on whether the task involved color or position adjustments. In color adjustment tasks, including changing image hue, applying fall colors, and adjusting color balance, many users preferred color wheels and color pickers for predictability and explorability, while preset buttons were preferred for efficiency, similar to continuous and discrete value adjustment tasks. However, for tasks requiring precise color adjustment, such as matching an object’s color to a given color, there was no clear consensus regarding predictability. Participants were evenly split among text fields, color wheels, color pickers, and preset buttons. The reasons for their preferences included “Since I know what my target is, just typing it gives me what I want” for the text field, “It is easy to visualize the color” for the color wheel, “The color picker would accurately allow me to pinpoint the target color green and not have a doubt that it is correct” for the color picker, and “The preset button convey closely what color the rocket will have” for the preset buttons. Regarding efficiency, most users preferred text fields, as they could quickly “enter the hex number and see the desired result”. For explorability, color wheels and color pickers were the most favored.
In position adjustment tasks, including placing a watermark or highlighting a specific area of an image, preset buttons and directly clicking on the image were preferred for predictability and efficiency, with reasons like “The presets give an indication of what to expect and wouldn’t confuse me in this task” and “Clicking is both intuitive and efficient and I know what I’$l$ get”. For explorability, users favored both clicking on the image since it was “Easy to just click in slightly different places as needed to get the area where I want it”, and using sliders because “You can totally visually discern where you want the watermarks and/or image to go to as it is continuous”.
These collected user preferences are incorporated to guide UI widget reasoning in our user evaluation of the LLM-driven UI widget generation system. We present the user evaluation details and how the crowdsourced user preferences affect reasoned widgets for user evaluation tasks in the following section.
7. User Evaluation
In this section, we introduce our user evaluation to assess our implemented system for the crowdsourced library-enhanced UI widget generation framework. We outline the research questions to evaluate in the study, describe the evaluation setup, and present the evaluation results.
7.1. Research Questions
We aim to study the following research questions in our user evaluation, focusing on the effectiveness and generalizability of the UI widget generation framework and the impact of integrating the crowdsourced UI widget preference library:
  • RQ1—Adaptivity of the framework to new, related tasks: Does the UI widget generation framework enable the generalization of crowdsourced library-enhanced widget reasoning and generation to perform tasks that differ from, but relate to, the original crowdsourcing tasks?
  • RQ2—Impact of the library on aligning generated widgets with user preferences: Can the crowdsourced UI widget preference library, which is integrated into the UI widget generation framework, help generated widgets align more closely with user preferences for predictability, efficiency, and explorability compared to those generated without a library?
  • RQ3—Effect of library size on enhancing LLM reasoning: Does the size of the crowdsourced UI widget preference library affect LLM widget reasoning effectiveness? Which size of the library provides the strongest support for LLM reasoning capabilities?
In the following subsections, we introduce the user evaluation setup designed to answer these research questions and analyze our evaluation results in alignment with them.
7.2. Platform and Participants
[…]
7.3. Evaluation Setup
7.3.1. Procedure
The user evaluation invited voluntary participants, who were first presented with a consent declaration at the beginning of the study. Once participants agreed to participate, they accessed the study interface via a URL linked to our server. Before starting the tasks, they were given a briefing that outlined the study requirements. Full details of the user evaluation briefing and instructions can be found in Appendix C. Following that, they worked on three tasks and answered six questions for each task to provide their preferences for LLM-generated widgets.
7.3.2. User Interface
Figure 7 shows the user interface for a task in the user evaluation—adjusting image tint. Each participant interacted with a Jupyter Notebook containing LLM-generated UI widgets to work on tasks distinct from those in the crowdsourcing study. The UI widgets were generated by the LLM alone and enhanced by our crowdsourced library. After interacting with all the offered widgets, participants were shown a preference selection panel where they chose their favored options for each task.
(a) Task description and LLM-generated UI widgets to perform the task.
(b) User preference selection panel for pairwise comparison. The panel contains six questions for each task, and we show one of them due to space limits.
Figure 7. User evaluation user interface for task: adjusting image tint (image_adjust_tint).
7.3.3. Tasks and Preference Aspects
The user study tasks differ from the crowdsourcing tasks but still fall within the same three categories, promoting either exploration or precision (recall Section 6.2.2). We provide an overview of these tasks and their categories below, with detailed descriptions available in Table 2. These tasks are designed to address RQ1. Note that two tasks belong to the design domain, i.e., design_align_text and design_position_logo, expanding beyond the image editing domain emphasized in the crowdsourcing study. Through these user evaluation tasks, we aim to test our framework’s generalizability to new tasks in both image editing and additional domains.
  • Continuous value adjustment: Tasks include adjusting image exposure, tint, and temperature.
  • Discrete value selection: Tasks include choosing predefined values for image exposure, tint, and temperature adjustment, text alignment, and logo positioning.
  • Tasks emphasizing visual feedback: Tasks include changing an image’s color to spring colors, aligning text with the margin, and placing a logo.
Table 2. User evaluation task names and descriptions.
Task Name Task Description
1 image_adjust_exposure Adjust the image exposure by both decreasing and increasing it. Find the exposure level that appears the most natural and visually pleasing to you.
2 image_adjust_tint Adjust the image tint to shift its color balance, creating both subtle and dramatic effects by adding a yellow or magenta hue.
3 image_adjust_temperature Adjust the image temperature to warm and cool tones. Experiment with different settings to achieve the most desired effect for you.
4 image_change_to_spring Transform the image to reflect vibrant spring colors, adding fresh, lively hues.
5 design_align_text Align the text “Poster Title” perfectly along one of the margins. Experiment with various positions to achieve a balanced and visually pleasing design.
6 design_position_logo Experiment with various placements for the logo to determine the most visually appealing position that enhances visibility and complements the overall design of the image.
The preference aspects: predictability, efficiency, and explorability, are drawn from the crowdsourcing study. We use the same aspects to evaluate whether user preferences collected from crowdsourcing can effectively lead to preference-aligned widget generation for tasks in the user evaluation. This setup addresses RQ2.
7.3.4. UI Widgets for the Evaluation Tasks
Figure 8 shows an example of LLM-reasoned widgets for task image_adjust_tint enhanced by three different sizes of the crowdsourced preference library: withlib10, withlib25, and withlib30, as well as without any library: withoutlib. Withlib10, withlib25, and withlib30 stand for the crowdsourced library with 10, 25, and 30 unique responses to each preference aspect. As introduced in Section 6.3.1, we obtain 30 responses to each aspect in total, i.e., withlib30, and use its subsets to construct withlib10 and withlib25.
In Figure 8, the percentages in the pie charts show the frequency of each type of UI widget reasoned by the LLM over 10 iterations. For instance, in the pie chart for efficiency reasoned by withoutlib (the rightmost chart in the second row), the preset buttons are reasoned 8 times (80%), while the slider is reasoned 2 times (20%). Reasoned widgets differing in multiple iterations are caused by the ad-hoc nature of LLMs, i.e., they do not always generate all possible widgets in a single pass, similar to the observations in previous work (Barke et al., 2023; Vaithilingam et al., 2024). This observation reveals varying consistencies of LLM-reasoned widgets with different library sizes. Thus, we leverage the reasoning frequency to indicate the recommendation strength for each type of widget by assigning each a score in our user evaluation. For example, the frequencies of 8 for the preset buttons and 2 for the slider are used as scores 8 and 2 for them, as seen in Figure 7(b) Option 2. Due to space limit, we provide the LLM-reasoned widgets with different library sizes for all six tasks in Appendx A.
As seen in Figure 8, LLM-reasoned UI widgets differ significantly with and without the crowdsourced libraries across the three preference aspects. For predictability, the color wheel is the most frequently reasoned widget when using the library, while the slider takes precedence in the absence of a library. In terms of efficiency, both withlib10 and withlib30 favor the color picker, while withlib25 most frequently reasons the color wheel, followed closely by the preset buttons and color picker. Without a library, preset buttons are the most commonly reasoned widgets. Regarding explorability, the color wheel is reasoned the most when libraries are utilized, whereas the color picker is the top choice without a library. Similar reasoning differences can be found in the reasoned widgets for other tasks in Appendx A. These findings prompt us to address RQ3: What is the optimal size of the crowdsourced library that maximizes LLM reasoning capabilities preferred by users with regard to the three preference aspects? To investigate this question, we conduct pairwise comparisons between libraries and collect user reasons for their choice.
Figure 8. LLM-reasoned UI widgets for image_adjust_tint with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
7.3.5. Pairwise Comparisons
We conduct pairwise comparisons for the reasoned widgets, their reasoned frequency, and the LLM-provided reasoning enhanced by withlib10, withlib25, withlib30, and withoutlib. By including widgets reasoned over multiple iterations, 10 in our study, we aim to include as many relevant widgets as possible in the evaluation, enabling a comprehensive comparison of LLM reasoning capabilities augmented by different sizes of the libraries. Participants need to select their preferences by comprehensively considering the widgets, their frequencies, where higher frequency indicates a stronger recommendation, and the reasoning, which explains why each widget is deemed appropriate for the task and preference aspect (recall Figure 7(b) for an example of pairwise preference selection). Each preference aspect in each task results in six pairwise comparisons, as outlined in Table 3.
Table 3. Pairwise comparisons of different crowdsourced library sizes for each preference aspect in each task. Pair orders are randomized in the user evaluation.
Library Sizes Compared
1 withlib10 vs. withlib25
2 withlib10 vs. withlib30
3 withlib10 vs. withoutlib
4 withlib25 vs. withlib30
5 withlib25 vs. withoutlib
6 withlib30 vs. withoutlib
In total, this gives us 6 pairs $\times$× 3 aspects $\times$× 6 tasks = 108 comparisons. To prevent user fatigue, each participant was assigned 18 pairs, which consisted of all six pairs for each aspect across three tasks. Specifically, the user study was divided into two task sets, each containing three tasks. Each participant provided preferences for all three aspects of their assigned task set. The tasks were arranged in all six possible order combinations, and the aspect order was balanced using the Latin Square method (Colbourn, 2010). See Table 4 for the detailed study setup. The six pairwise comparisons for each aspect of each task were presented in random order. To ensure robust statistical results, each pairwise comparison was tested by 12 users, resulting in 6 pairs $\times$× 12 users = 72 total participants.
Table 4. User evaluation setup. The six user evaluation tasks are divided into two task sets. In each set, we arrange all six order combinations of tasks and aspects and randomize orders of pairwise comparisons.
Task Set 1 Task Set 2
Order Task Name Aspect Order Task Name Aspect
6 $\times$× image_adjust_exposure Predictability 6 $\times$× image_adjust_tint Predictability
image_adjust_temperature Efficiency image_change_to_spring Efficiency
design_align_text Explorability design_position_logo Explorability
6 $\times$× image_adjust_exposure Efficiency 6 $\times$× image_adjust_tint Efficiency
image_adjust_temperature Explorability image_change_to_spring Explorability
design_align_text Predictability design_position_logo Predictability
6 $\times$× image_adjust_exposure Explorability 6 $\times$× image_adjust_tint Explorability
image_adjust_temperature Predictability image_change_to_spring Predictability
design_align_text Efficiency design_position_logo Efficiency
7.4. Results
7.4.1. Results of Each Aspect across All Tasks
We present the user evaluation results of each preference aspect across all tasks in Figure 9. As can be seen, for predictability and explorability, users preferred the LLM-reasoned widgets with libraries over those without any library. In terms of efficiency, LLM-reasoned widgets with libraries were slightly more favored than those without. When the libraries were incorporated, withlib30 was more preferred than withlib10 and withlib25 for predictability and explorability, while for efficiency, withlib25 and withlib30 were almost equally favored and more preferred than withlib10.
Figure 9. Total user preference selection counts of three preference aspects across all tasks.
7.4.2. Results of Each Aspect in Each Task
Next, we examine user preferences for LLM-reasoned widgets per aspect and task. Figures 10 and 11 depict results of the task sets 1 and 2. As observed, user preferences show varying trends depending on the task.
Figure 10. User preference selection counts of tasks in task set 1.
Figure 11. User preference selection counts of tasks in task set 2.
In task set 1, for image_adjust_exposure (see Figure 13 for LLM-reasoned widgets), widgets reasoned with libraries were more preferred than those without a library for both predictability and explorability, with withlib25 and withlib30 being nearly equally favored. For efficiency, withlib25 and withlib30 were preferred over withoutlib, and withoutlib was favored over withlib10. When LLM-reasoned widget types were similar by different libraries, such as 60% preset buttons and 40% slider by withlib25, and 70% preset buttons and 30% slider by withlib30, participants’ feedback suggested that the slight score difference and LLM-provided reasoning helped guide their preference selection. For instance, participants chose withlib30 due to “good reasonings for any user which would prefer an easier time with Preset Buttons of higher score”.
For image_adjust_temperature (see Figure 15 for LLM-reasoned widgets), withoutlib was preferred for predictability and almost equally favored as withlib25 and withlib30 for efficiency. Participants who preferred sliders reasoned by withoutlib over the color wheel and color picker reasoned with the libraries highlighted the slider’s predictability. For example, they commented, “the slider is the most instantly understandable widget… the color wheel and color picker require a bit of deduction from the user on how to utilize them properly”. They also remarked on sliders’ sensitivity to value changes, such as “I like the sensibility of sliders in general, and I feel they are a very efficient means of getting a good result by eye, just by dragging your mouse left or right, so this is the best”. However, widgets reasoned with libraries were favored for explorability. For instance, participants preferred the color picker reasoned by withlib10 because “you are given 3 RGB values each ranging from 0 to 255, giving you 255^3 possibilities”, as opposed to the slider reasoned by withoutlib since “you are only given 100 options as the slider moves in intervals of 0.01”.
With regard to design_align_text (see Figure 17 for LLM-reasoned widgets), widgets reasoned with libraries were again more favored for predictability, and trends similar to image_adjust_temperature were observed for efficiency and explorability.
In task set 2, for image_adjust_tint (see Figure 8 or 14 for LLM-reasoned widgets), widgets reasoned with libraries remained more favored for predictability compared to those without a library. For efficiency, withoutlib was the most preferred option, indicating that the preset buttons were favored over the color picker and color wheel reasoned with libraries. Regarding explorability, withlib10, withlib25, and withoutlib were nearly equally favored, showing that participants preferred the color wheel and color picker for exploration over sliders reasoned by withlib30.
In terms of image_change_to_spring (see Figure 16 for LLM-reasoned widgets), withlib25 and withlib30 were almost equally preferred, both more favored than withlib10 and withoutlib for predictability. In terms of efficiency, withlib10 and withlib25 were preferred over withlib30 and withoutlib. For explorability, widgets reasoned with libraries were consistently more favored than those without any library.
Regarding design_position_logo (see Figure 18 for LLM-reasoned widgets), widgets reasoned with libraries were more favored across all three aspects. Withlib10 was the most preferred for predictability and efficiency, while withlib30 was the most favored for explorability. For predictability and efficiency, withlib10 reasoned similar widgets (click on image) to those reasoned by other libraries, but it was preferred over withlib25 and withlib30 largely due to its reasoning. For instance, participants commented, “Option 1 [withlib10] made more sense because it said how it takes minimal effort, instant feedback, and it’s quick and precise. Whereas option 2 [withlib30] said it’s just simple and takes minimal effort”.
7.4.3. Reflection on Research Questions
For RQ1, the results reveal that the user preferences collected from crowdsourcing can effectively generalize to tasks beyond the original crowdsourcing tasks by enhancing the LLM’s reasoning capabilities. Broadly speaking, across all user evaluation tasks, widgets reasoned with the library are largely favored over those reasoned without a library. This highlights the adaptability and relevance of user preferences captured through crowdsourcing, suggesting that these insights can serve as a foundation to inform widget reasoning across a range of related tasks. While the overall trend favors library-informed reasoning, some exceptions emerge for specific tasks. For instance, for predictability in image_adjust_temperature and both efficiency and explorability in image_adjust_tint, widgets reasoned without the library are preferred. These results suggest that certain tasks may require additional consideration in the reasoning process.
For RQ2, the results demonstrate that incorporating the crowdsourced library enhances the ability of the LLM to reason widgets that align with user preferences regarding predictability, efficiency, and explorability. In terms of predictability, users’ preferences captured from the crowdsourcing study lean towards preset buttons for predefined value selection, color wheels, and color pickers for color adjustments, and click on image for position changes (recall Section 6.3.2). These preferences have directly influenced the widgets reasoned for the user evaluation tasks, where widgets reasoned with the library adhered more closely to these favored widget types. In contrast, when the library is not used, the LLM reasons sliders for most tasks regardless of the task categories. For efficiency, similar patterns emerge—preset buttons, color wheels, color pickers, and click on image are repeatedly preferred for their ease and speed in achieving task goals. The LLM, when equipped with the crowdsourced library, is able to reason toward these efficient widgets, whereas, without the library, the reasoning leans more heavily on sliders and preset buttons for most tasks without too much differentiation for task categories. Regarding explorability, the user preference trends show that sliders are favored for continuous value adjustments, color wheels are preferred for color changes, and click on image is liked for position adjustments. Without the library, however, the LLM tends to reason sliders to allow exploration for continuous value adjustments, color pickers for color changes, and sliders and text fields for position adjustments. The incorporation of the crowdsourced library enables the LLM to capture the need for exploration in tasks like color change and position adjustment by effectively aligning the widget choices with user preferences.
For RQ3, we observe that withlib30, containing the most crowdsourced data, emerges as the most preferred library across participants, particularly for predictability and explorability, while performing almost equally with withlib25 in terms of efficiency. This pattern, supported by the distribution of user preferences across all tasks, as shown in Figure 12, highlights withlib30 as the most reliable for all three aspects due to the richness of its crowdsourced data. This consistency of user preferences underscores the value of extensive, diverse data in reasoning well-aligned widget options for users. On the other hand, withoutlib shows more significant variability in user preferences, particularly in predictability and efficiency, suggesting that when no library is used, LLM reasoning lacks the consistency that is otherwise afforded by crowdsourced data. While withlib10 and withlib25 are moderately favored for predictability and efficiency, they do not perform as well in explorability as withlib30, further emphasizing the benefit of larger datasets in supporting widget reasoning for tasks requiring exploration. However, examining each task individually reveals significant variation in preferred libraries for different preference aspects. This indicates that while larger libraries lead to consistent preference-aligned widget reasoning across multiple tasks in general, for individual tasks, the utility of LLM-reasoned widgets for specific preference aspects may need additional evaluation.
Figure 12. Distribution of user preference selection counts of three preference aspects across all tasks.
8. Discussion
Based on the analysis of user evaluation results, this section discusses the implications of these results from three perspectives: the balance between task-generic and task-specific LLM widget generation, the alignment with user preferences for user-centered UI design, and the use of user preference data to create UIs that adapt to personalized use cases and evolve over time.
8.1. Generalizability of User Preference Data to Support Various Tasks
The user evaluation findings suggest that crowdsourced UI widget preference libraries provide an effective foundation to improve LLM reasoning across tasks beyond the original scope of crowdsourcing. From the tasks tested in the user evaluation, we have shown the task generalizability not only in the image editing domain but also in the design domain (recall tasks 5 and 6 in Table 2). This generalization ability can reduce the need for task-specific widget customization using automatic LLM widget generation, streamlining the UI design process. To further enhance this generalizability, the scope of crowdsourced tasks can be expanded by incorporating a more comprehensive array of task categories beyond continuous value adjustment, discrete value selection, and color and position changes. In addition, we can deepen the exploration within specific task categories to involve more complex aspects of data manipulation. For instance, our study emphasized tasks that required exploration and precision within the three task categories, and this can be extended to tasks on content editing or value adjustment that emphasize simplicity, speed, or flexibility. Also, while we focused on single-value adjustment, this setup can be extended to multi-dimensional data manipulation or compound parameter adjustments. By carefully curating the tasks for crowdsourcing, we envision that LLMs can be equipped with crowdsourced insights to reason UI widgets in both task-generic and task-specific contexts, adapting to diverse user needs and task scenarios.
Moreover, combining task-generic and task-specific UI widget reasoning could provide a solution for cases where generalized user preferences alone are insufficient. In our user evaluation, preference aspects of a few tasks showed that user preferences were better supported without the library (recall Sections 7.4.2 and 7.4.3). These findings highlight that crowdsourced preferences do not universally apply to all tasks, and some require more refined, context-sensitive reasoning. To address this, an adaptive approach that integrates task-specific customization with the general crowdsourced library can potentially better accommodate unique user interaction demands to ensure the LLM delivers more tailored and effective UI widget suggestions across a wide range of tasks. For example, for the efficiency aspect of image_adjust_tint, users showed a preference for preset buttons generated by withoutlib over the color wheel and color picker reasoned with the library. Incorporating this insight into the LLM reasoning process by assigning higher weights to preset buttons for tasks involving image tint adjustments and related operations can lead to more preference-aligned UI generation than directly using the library. This approach leverages the flexibility of LLM-based widget generation by taking updated user insights to adapt to both generic and specific task requirements.
8.2. Towards User-Centered UI Widget Selection and Design
The crowdsourcing results reveal significant variation in user preferences for UI widgets based on different data manipulation needs, emphasizing the importance of a user-centered approach to UI widget selection. This approach contrasts with the traditional designer-dominant method, where UI widgets are often pre-selected by software designers. For instance, tools like Photoshop provide a slider and a text field for adjusting image lightness, saturation, and hue, which offer a one-size-fits-all solution. However, our crowdsourced data shows that users prefer distinct widgets for various data manipulation requirements, such as preset buttons for efficient adjustment of lightness and saturation and a color wheel to explore multiple hue options (recall Section 6.3.2). This divergence reveals a gap between user-preferred and designer-selected UI widgets. This gap highlights that static, predefined UI widgets in most software may not always meet the diverse needs of users. LLM-supported UI generation offers a promising solution to bridge this gap by generating dynamic widgets. With the crowdsourced preference library, we have extended the existing LLMs’ ability to reuse user-preferred widgets based on user preference insights and, therefore, generate preference-aligned widgets for a broad range of tasks. This framework can be used to enable the creation of user-centered interfaces that align with the preferences collected from users to augment the traditional UI widget selection method that heavily relies on designers.
Furthermore, the user-centered UI design approach can be expanded to accommodate a broader spectrum of preference aspects and cater to more diverse user groups. Our study focuses on three aspects of user preferences, demonstrating the prototype’s effectiveness in guiding LLM reasoning, but future research could incorporate additional dimensions of UI design. For example, prior work on UI design principles (McKay, 2013; Lidwell et al., 2010; Chen et al., 2021), which covers a number of guidelines, can be used to inform crowdsourcing efforts. Among them, accessibility is a particularly important dimension that can be addressed by collecting user preferences from individuals with disabilities. Crowdsourced data from these users could inform the design of accessible UI widgets, ensuring that interfaces meet their specific needs. This method would not only aid LLMs in generating accessible widgets that benefit users but also provide valuable insights for software designers. For example, LLMs can be equipped with crowdsourced data to suggest widget sizes optimized for low-vision users or alternative interaction techniques for those with motor impairments. They can also be prompted to offer reasonings for their suggestions, which could serve as references for design rationale to assist designers. Thus, the LLM-reasoned widgets can potentially augment the ideas proposed in prior research, such as SUPPLE (Gajos and Weld, 2004) and SUPPLE++ (Gajos et al., 2007), to develop dynamic UIs that accommodate motor and vision capabilities.
8.3. Opportunities for Adaptive UIs
The comparison between different library sizes reveals that larger libraries tend to produce more consistent preference-aligned widget reasoning across various tasks; meanwhile, it highlights the potential of smaller libraries for cost-effective and adaptive UI design. Our study shows that even with a library of just 10 user responses, i.e., withlib10, key insights into user preferences can be effectively captured. In some tasks and aspects, such as the predictability of adjusting image exposure and tint, the efficiency of changing colors to spring tones and positioning a logo, and the explorability of aligning text and adjusting tint, smaller libraries performed as well or even better than larger ones. Small-scale libraries are quick to collect and easy to evolve, providing efficient UI design suggestions for designers to reference in the design process. Even at the extreme, i.e., using no library at all, LLMs can reason UI widgets that, in some instances, outperformed those based on larger libraries. This demonstrates that LLMs can aid in UI widget reasoning and selection without requiring extensive human input, which offers designers a low-cost way to access user preferences during UI design. Such adaptability of small libraries can significantly reduce the time and resources required to gather large datasets while still yielding meaningful user-centered insights.
The adaptability enabled by small-scale libraries for UI design can be extended to deeper UI personalization, which reflects individual preferences and supports long-term UI evolution. By collecting a few samples from individual users, these libraries can be tailored to represent personalized preferences for UI widget generation. Such personalization can allow users to build workflows that align with their specific needs. Moreover, personalized libraries have the potential to evolve alongside users as their skills and expertise grow. For example, the LLM could initially generate simple, beginner-friendly widgets and, over time, adapt to offer more advanced, feature-rich widgets as the user becomes more proficient by using the user’s updated personal library it uses for widget reasoning. This dynamic evolution enables users to progressively engage with more complex tools to support an adaptive and long-term learning curve. This kind of UI evolution not only streamlines immediate tasks but also fosters continuous improvement in tool use, helping users to gradually master more sophisticated functions without overwhelming them at the start.
9. Limitations and Future Work
The first limitation lies in the range of preferences covered during crowdsourcing. Our study focused on three key aspects, but there are other significant UI design principles, such as accessibility, cognitive load, and adaptability (Lidwell et al., 2010), that were not explicitly addressed here. However, our framework is highly extendable and can be enhanced by incorporating a broader set of preference dimensions with customizable aspects to account for user preferences from different perspectives.
Next, we mainly considered software-based UI widget generation in this work. While these widgets suited our tasks, other potential options, like gesture-based or tangible widgets, were not included. This may limit our understanding of how different widgets, like those supporting spatial interactions in XR (Aponte et al., 2024) or physical interactions in tangible interfaces (Campos Zamora et al., 2024), could meet specific task needs or user preferences. Expanding the widget types in future research could offer comprehensive insights into user preference across a broader spectrum of tasks and interaction styles.
Lastly, our system implementation is currently focused on UI generation within the image editing domain. In the future, we aim to collect data from multiple domains and various tasks to build a more extensive user preference library. This would allow us to evaluate how a cross-domain library can improve widget generation, both for in-domain tasks and for tasks outside the collected domains. Such a system would also introduce challenges, such as how to effectively manage large datasets and ensure the LLM fully leverages all the available knowledge.
10. Conclusion
In this paper, we introduced a UI widget preference library collected from a crowdsourcing study and a framework for LLM-driven widget generation enhanced by this library. Our evaluation of the framework demonstrated its ability to extend widget generation across diverse but related tasks, align generated widgets with user preferences captured through crowdsourcing, and use large libraries for consistent, preference-aligned widget reasoning while leveraging smaller libraries for cost-effective preference capture. We further discussed these results to explore the extendability of crowdsourcing tasks, preference aspects, and library sizes in supporting UI generation tailored to task-generic and task-specific use cases, diverse user preferences and needs, and adaptive workflows. We envision our framework to benefit both users and designers by simplifying the development of user-centered interfaces that enhance task efficiency and user satisfaction.
Appendix A LLM-Reasoned UI Widgets for User Evaluation Tasks
Figures 13 to 18 show the LLM-reasoned UI widgets with crowdsourced libraries and without any library for the user evaluation tasks.
Figure 13. LLM-reasoned UI widgets for image_adjust_exposure with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
Figure 14. LLM-reasoned UI widgets for image_adjust_tint with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
Figure 15. LLM-reasoned UI widgets for image_adjust_temperature with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
Figure 16. LLM-reasoned UI widgets for image_change_to_spring with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
Figure 17. LLM-reasoned UI widgets for design_align_text with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
Figure 18. LLM-reasoned UI widgets for design_position_logo with three sizes of the crowdsourced libraries and without any library. Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
Appendix B Prompts for LLM UI Widget Reasoning and Code Generation
B.1. Prompts for UI Widget Reasoning without Crowdsourced UI Preference Library
Reason UI widget type for the user task. You should follow the steps below for the reasoning.
First, please take the definitions below:
- Predictability: allows users to obtain results with no surprises or doesnt require users to deduce how to perform the interaction.
- Efficiency: allows users to perform tasks with a minimum amount of effort.
- Explorability: allows users to explore multiple possibilities and perform functions with high flexibility.
Second, reason the most proper UI widget for predictability, efficiency, and explorability.
- The UI widget you reason must from Slider, Preset Buttons, Dropdown, Radio Buttons, Text Field, Color Wheel, Color Picker, Click on Image.
Third, based on your reasoning, write down UI widgets for predictability, efficiency, and explorability in JSON format.
Refer to the example below to provide your response.
- Replace the placeholders marked by <> with your response. Do not include <> in your response.
- You must keep all the existing information from the example and only replace placeholders.
- The response must be in JSON format.
"reasoning": {‘‘‘reasoning
"task analysis": "<your reasoning>"
"predictability_reasoning": {
"<UI widget type>": "<your reasoning>"
"efficiency_reasoning": {
"<UI widget type>": "<your reasoning>"
"explorability_reasoning": {
"<UI widget type>": "<your reasoning>"
"widget": {‘‘‘widget
"<task_name>": {
"predictability": "<UI widget type>",
"efficiency": "<UI widget type>",
"explorability": "<UI widget type>"
B.2. Prompts for UI Widget Reasoning with Crowdsourced UI Preference Library
Based on the crowdsourced UI widget preference library, reason UI widget type for the user task. You should follow the steps below for the reasoning.
First, please take the definitions below:
- Predictability: allows users to obtain results with no surprises or doesnt require users to deduce how to perform the interaction.
- Efficiency: allows users to perform tasks with a minimum amount of effort.
- Explorability: allows users to explore multiple possibilities and perform functions with high flexibility.
Second, information on the crowdsourced UI widget preference library is in the prompt:
- Crowdsourced UI widget preference library task description: detailed descriptions of all the tasks.
- Crowdsourced UI widget preference library widget frequency: the frequency of user-preferred widgets. Large numbers mean the corresponding widget is preferred by more people.
- Crowdsourced UI widget preference library widget reasons: the reasons for user-preferred widgets.
Third, search for the most relevant tasks from the crowdsourced UI widget preference library.
- You can compare the given task and the tasks names in the library, and refer to their task descriptions to help you to find the relevant tasks.
- Additional task information:
- "image_adjust_lightness", "image_adjust_saturation", and "image_adjust_hue" represent tasks that allow continuous value adjustment and discrete value selection, and "image_adjust_hue" is also related to color adjustment.
- "image_adjust_fall_color", "image_color_match", and "image_adjust_color_balance" are related to color adjustment.
- "image_place_watermark" and "image_place_vignette" are related to object positioning and discrete value selection.
Fourth, reason the most proper UI widget for predictability, efficiency, and explorability.
- Your reasoning should be based on the relevant tasks you found in the library.
- After you find the relevant tasks, refer to the content of "Predictability", "Efficiency", or "Explorability" in the widget frequency and widget reasons.
- You must refer to the widgets of high frequencies of the relevant tasks in widget frequency.
- You must refer to widget reasons to help your reasoning.
- The UI widget you reason must come from the given library.
Lastly, based on your reasoning, write down UI widgets for predictability, efficiency, and explorability in JSON format.
Refer to the example below to provide your response.
- Replace the placeholders marked by <> with your response. Do not include <> in your response.
- You must keep all the existing information from the example and only replace placeholders.
- The response must be in JSON format.
"reasoning": {‘‘‘reasoning
"relevant tasks from the library": "<your reasoning>"
"predictability_reasoning": {
"<UI widget type>": "<your reasoning>"
"efficiency_reasoning": {
"<UI widget type>": "<your reasoning>"
"explorability_reasoning": {
"<UI widget type>": "<your reasoning>"
"widget": {‘‘‘widget
"<task_name>": {
"predictability": "<UI widget type>",
"efficiency": "<UI widget type>",
"explorability": "<UI widget type>"
B.3. UI Preference Library Appended to the UI Widget Reasoning Prompt
Crowdsourced UI widget preference library task description:
"image_adjust_lightness": "Adjust the lightness of the image.",
"image_adjust_saturation": "Adjust the saturation of the image.",
"image_adjust_hue": "Adjust the hue of the image.",
"image_adjust_fall_color": "Change the image to fall colors.",
"image_color_match": "Match the color of the image to a given target color.",
"image_adjust_color_balance": "Adjust the color balance of the image by adjusting the RGB channels.",
"image_place_watermark": "Place a watermark on the image to balance between visibility and subtlety.",
"image_place_vignette": "Darken the background area using the vignette effect except for a certain image area."
Crowdsourced UI widget preference library widget frequency:
"image_adjust_lightness": {
"Predictability": {
"Slider": 1,
"Preset Buttons": 4
"Efficiency": {
"Slider": 3,
"Text Field": 1,
"Radio Buttons": 1
"Explorability": {
"Slider": 5
"image_adjust_saturation": {
"Predictability": {
"Slider": 2,
"Preset Buttons": 3
"Efficiency": {
"Slider": 1,
"Preset Buttons": 4
"Explorability": {
"Text Field": 1,
"Slider": 4
"image_adjust_hue": {
"Predictability": {
"Color Wheel": 4,
"Preset Buttons": 1
"Efficiency": {
"Slider": 1,
"Color Wheel": 1,
"Preset Buttons": 2,
"Radio Buttons": 1
"Explorability": {
"Color Wheel": 4,
"Slider": 1
"image_adjust_fall_color": {
"Predictability": {
"Color Wheel": 4,
"Preset Buttons": 1
"Efficiency": {
"Preset Buttons": 2,
"Slider": 1,
"Radio Buttons": 1,
"Color Wheel": 1
"Explorability": {
"Color Wheel": 4,
"Slider": 1
"image_color_match": {
"Predictability": {
"Text Field": 2,
"Color Picker": 3
"Efficiency": {
"Text Field": 1,
"Radio Buttons": 1,
"Color Picker": 1,
"Preset Buttons": 2
"Explorability": {
"Color Wheel": 4,
"Color Picker": 1
"image_adjust_color_balance": {
"Predictability": {
"Color Picker": 4,
"Slider": 1
"Efficiency": {
"Color Picker": 3,
"Text Field": 1,
"Slider": 1
"Explorability": {
"Slider": 3,
"Color Picker": 2
"image_place_watermark": {
"Predictability": {
"Click on Image": 3,
"Preset Buttons": 1,
"Dropdown": 1
"Efficiency": {
"Click on Image": 2,
"Preset Buttons": 3
"Explorability": {
"Preset Buttons": 1,
"Click on Image": 2,
"Slider": 2
"image_place_vignette": {
"Predictability": {
"Preset Buttons": 2,
"Click on Image": 2,
"Slider": 1
"Efficiency": {
"Click on Image": 3,
"Slider": 2
"Explorability": {
"Click on Image": 2,
"Preset Buttons": 1,
"Text Field": 1,
"Slider": 1
Crowdsourced UI widget preference library widget reasons:
"image_adjust_lightness": {
"Predictability": {
"Slider": [
"Its easy to understand that higher values will produce lighter images and viceversa"
"Preset Buttons": [
"The preset buttons background allows for users to figure out the lightness relatively quickly, and when clicking on the buttons, the button updates the lightness to how it looks on the buttons, allowing for users to need minimal deduction to understand what this button does.",
"It is like to have an example of what you getting as a result",
"The numbers show how each setting may look.",
"the preset buttons provided an easy adjustment to the picture."
"Efficiency": {
"Slider": [
"I can move it to the area I like the most",
"Easier to adjust and more intuitive ",
"the slider felt like the most efficient method to adjust the exposure. This method being something I am familiar with changing in Adobe programs."
"Text Field": [
"If a user were more experienced in photo editing, they can quickly adjust the lightness precisely to what they want it to be by simply typing in the value, making it quick and easy."
"Radio Buttons": [
"The values are displayed clearly and you just click one."
"Explorability": {
"Slider": [
"Fast and contains options to the valid range so I dont have to figure that out.",
"It allows a greater range with more fluid motion, it provides the result on the spot, and it feels natural. Can select more ranges from a wide variety to see different types of results.",
"Allows for a more precise lighting value, and quicker",
"Theres a bigger range of options",
"I find there is a lot of sensitivity in the slider, which has greater impact."
"image_adjust_saturation": {
"Predictability": {
"Slider": [
"Saturation is usually set through a slider in image editing software",
"using the slider was the easiest method to adjust the saturation. "
"Preset Buttons": [
"The preset buttons include a preview of how the saturation would look like. For example, -1.0 is fully gray while 1.0 is a bright rainbow. This allows for users who may not know what saturation means to understand what each button does, it changes how \"colorful\" the image is.",
"It is easy to see the level of saturation that you want",
"It shows what the colors are like and then I go from there."
"Efficiency": {
"Slider": [
"Fast and easy to set it to a value I like."
"Preset Buttons": [
"The ability to view gradients and brightness on the preset buttons really gave me an advantage. I had an idea of the color spectrums brightness and what to expect, while the other options was almost like choosing in the dark and waiting for a surprise.",
"Easier to use, and provides a glimpse into the saturation levels that radio buttons do not.",
"You can just click the button and see the result assuming its a simple task like this one to just make colors pop",
"This makes editing easier, as the values are predetermined."
"Explorability": {
"Text Field": [
"I can enter specific values and see their results"
"Slider": [
"The slider allows users to slowly go through each increment and explore what effect they have on the image. They can go anywhere within the range to understand the effect of saturation on the image and precisely pinpoint which exact value of saturation they want with a slider.",
"It is intuitive to increase the level of saturation just with the slider bar",
"It discretely lists all possible colors and lets me just pick one while I scroll.",
"the slider also contributed to the explorability."
"image_adjust_hue": {
"Predictability": {
"Color Wheel": [
"Colors dont feel like they should map to an order, and the color wheel is the only option that doesnt try to order the colors.",
"For selecting the hue in this widget the color wheel was the most efficient choice in my opinion. It covers all the colors and gives me a clear area to select, the others I dont have much of guidance until I chose the options and see the results.",
"Theres a lot more options and you know exactly what youre getting when you press an option.",
"The color wheel provides more flexibility with color hues than the preset values."
"Preset Buttons": [
"The backgrounds of the preset buttons allow the user to already have a baseline as to what that hue value would then look like"
"Efficiency": {
"Slider": [
"Im more familiar with using a slider to change the hue of an image"
"Color Wheel": [
"If the task were to match the image to the same hue as another image, the user could look at the reference image and quickly pinpoint it to which hue it matches closely to on the color wheel, without needing to explore additional options."
"Preset Buttons": [
"You obtain the colour that you choose",
"the preset buttons provided the most efficient way to change the hue."
"Radio Buttons": [
"It gives a simple numeric representation which I can just click on."
"Explorability": {
"Color Wheel": [
"I can select different colours and check the results",
"There are more colour options ",
"The color wheel shows how each color hue looks so I can see all combinations from the outside.",
"the color wheel provided the most options to adjust the hue and explore other possibilities."
"Slider": [
"The slider allows for users to see the exact hue value as opposed to the color wheel, and see what even a small 0.01 increment in hue can do for an image. A slider for this task allows for more exploration as users scrub through the slider and see the different hue changes."
"image_adjust_fall_color": {
"Predictability": {
"Color Wheel": [
"Like before, colors dont have an order, so the wheel lets me know what I will get before clicking it.",
"By being able to select a color on the wheel and it produce the widget result with that color I will say its the most appropriate option for predictability. I wouldnt have to deduce much knowing the style of orange or yellow I chose is what will be applied.",
"Color wheel makes it easier to pick the colour you want, allowing for a more direct choice as you see the colours at once.",
"I find it easer to navigate the wheel due to be able to seeing the colours beforehand."
"Preset Buttons": [
"Simple interface that has preset values and you know what to expecct"
"Efficiency": {
"Preset Buttons": [
"I can quickly choose an option",
"You can choose exactly the colour you want for the picture just with one click"
"Slider": [
"For this specific task, the slider is the most efficient option, because the other preset options do not create a autumnal atmosphere. The preset buttons only allow for red, where autumn is more orange or brown. The slider allowed for quickly scrubbing through each value to see what was most warm and orange, and then fine tuning the smaller 0.01 values for the exact atmosphere vibes."
"Radio Buttons": [
"You have few choices and can just click a simple choice."
"Color Wheel": [
"The color wheel was the most efficient way to perform the task. I was able to click on which color I thought I would like."
"Explorability": {
"Color Wheel": [
"Its easier to test different possibilities and check the results",
"The color wheel allows the user to explore different atmospheres for the image without having to figure out what 0-1.0 means in this context. The color wheel allows for the overall images atmosphere to be adjusted and explored in different seasonal contexts like pink for spring, and blue for winter.",
"It lets you see the whole color spectrum and choose a particular hue that is right for the situation.",
"The color wheel also offered more visual help to identify the colors and make the task more accessible."
"Slider": [
"As you can choose different values and see how it is affecting on the picture"
"image_color_match": {
"Predictability": {
"Text Field": [
"Since I know what my target it, just typing it gives me what I want.",
"Can enter the hex number and see the desired result"
"Color Picker": [
"The color picker would accurately allow me to pinpoint the target color green and not have a doubt that it is correct.",
"As you know what colour you are picking there are no surprises",
"You can match the color visually before selecting it"
"Efficiency": {
"Text Field": [
"Being able to directly type what I want is the faster than trying to use the slider or color picker or wheel."
"Radio Buttons": [
"It would be the easiest way to click the options while producing the results under minimal effort."
"Color Picker": [
"Allows you to directly choose the target colour, with minimal effort/attempts"
"Preset Buttons": [
"Simple interface with preset options. Little to no effort required",
"This is much quicker to select, as the colors are preset"
"Explorability": {
"Color Wheel": [
"Easy to try out options by just clicking.",
"I feel like the wheel would have the most amount of options due to the color palette available to choose from.",
"Allows for many options to be explored quicker than any of the other options",
"This allows you to experiment more with the colors"
"Color Picker": [
"More explorability options and plenty of options to choose from."
"image_adjust_color_balance": {
"Predictability": {
"Color Picker": [
"Im familiar with using a colour picker to change the colours",
"The color picker widget adjusts the overall images color harmony to the picked color quickly, and does not make the process confusing for the user to figure out how much red, green, and blue they need for the image to turn out a certain color.",
"You need only to ckick the colour that you want to ",
"The widget just lets you see a color and overlay it on the image."
"Slider": [
"I had no surprises using the slider to adjust the color balance."
"Efficiency": {
"Color Picker": [
"It feels more intuitive",
"It is easy to choose a colour, however there are less options than for the slider",
"You can see what colors you want and just go with them without altering numbers."
"Text Field": [
"For more experienced users, the text field allows for them to efficiently type out the exact values they want to adjust without additional interaction. If the image needed just a little more blue, they could add 0.08 to it quickly, instead of fiddling with the slider."
"Slider": [
"The slider allowed me to perform the task with minimum effort. The slider felt more natural to use."
"Explorability": {
"Slider": [
"Fastest way to just try different options.",
"With multiple colors, the sliders makes it the most efficient and easy to use. It allows me to switch from colors easily and change their RGB values easily.",
"Is the quickest out of all the options to explore many values other than the colour picker, however, the colour picker makes it harder to change individual colour channels."
"Color Picker": [
"Allows you click into colors and enter manually the RGB values.",
"This allows users to explore more with the color settings"
"image_place_watermark": {
"Predictability": {
"Click on Image": [
"Clicking is both intuitive and I know what Ill get. It did take two clicks at least, because I thought the text would be centered on the click, but how it worked was quickly apparent so I could then adjust things as I wanted.",
"As the user directly clicks on the image, they know exactly where the watermark will be placed.",
"Text is there and you can just click on image and it will go exactly where you pressed."
"Preset Buttons": [
"The presets give an indication of what to expect and wouldnt confuse me in this task. I could choose the proper preset and its identifiable where the water mark would go."
"Dropdown": [
"It is easier to understand where the cursor will be positioned"
"Efficiency": {
"Click on Image": [
"I can simply click where I want the watermark to be",
"It is easier to choose the right position"
"Preset Buttons": [
"The typical locations a watermark would be placed in are listed in the preset buttons widget. In terms of efficiency, a user can quickly click on which location they want the watermark to appear and be done. This widget is also efficient in allowing the user to click through each option and explore where they want the watermark to be located quickly.",
"They just give you a few options set out for you to pick from.",
"The preset buttons are the most efficient way of adjusting the text."
"Explorability": {
"Preset Buttons": [
"I can test different variations easily"
"Click on Image": [
"If a user wants more control over where their watermark is placed, the Click on image widget option is the best for this, because the user can click on exactly where they want the watermark to be, then fine tune it by clicking slightly, ultimately adjusting their watermark to their preferred location.",
"You can totally visually discern where you want the watermarks and/or image to go to as it is continuous."
"Slider": [
"You can have more control by watching the numbers ",
"The slider made the explorability more accessible to adjust the text."
"image_place_vignette": {
"Predictability": {
"Preset Buttons": [
"Its easier to understand where the circle will move to",
"The preset buttons for this task allows for a user to quickly move the circle to common positions without any surprises. Users less familiar with photo editing can simply click on a button and know what it should be doing, and this widget does just that."
"Click on Image": [
"Very intuitive ",
"I just click on a location and it goes where I want."
"Slider": [
"I feel that the slider came with no surprises. It was the easiest way to adjust the position of the circle."
"Efficiency": {
"Click on Image": [
"The only option that really works well.",
"Lets the user quickly determine where the circle will be placed",
"You can point to exactly you need t be"
"Slider": [
"The easiest method would be the slider because it will produce the results with the least amount of effort. Its not the most powerful method but it is still efficient to a degree.",
"It was the easiest to get the best outcome without taking a long time."
"Explorability": {
"Click on Image": [
"Easy to just click in slightly different places as needed to get the circle where I want it.",
"As it is the quickest method it also allows the user to check many positions more quickly than the other options."
"Preset Buttons": [
"The presets give me the most ranges with options to choose from that are already covering a wide variety. Its flexible and also easy to use."
"Text Field": [
"Allows you enter exact co-ordinates giving more flexibility."
"Slider": [
"This provides more flexibility with the positioning"
B.4. Prompts for UI Widget Code Generation
Generate code for the UI widgets to perform the specified task. You should follow the steps below for the coding.
First, the UI widget code you provide must allow users to perform the task specified in the content of User task.
Second, the UI widget code you provide must follow the examples offered below for the implementation of Slider, Dropdown, Radio Buttons, Text Field, Preset Buttons, Color Wheel, Color Picker, Click on Image.
- The UI widget types are in widget_type.
- Find the example code for the relevant UI widget types.
- When coding, you must only change the task to the specified task, implement the specified widgets, and keep the remaining code format the same as the example code.
Third, write your response in JSON format following the example responses below.
"task": "Adjust image hue",
"widget_type": "Slider, Dropdown, Radio Buttons, Text Field, Preset Buttons, Color Wheel, Color Picker",
"widget_code": {‘‘‘python
import numpy as np
import matplotlib.pyplot as plt
import ipywidgets as widgets
from IPython.display import display, clear_output
from PIL import Image
from skimage import data, img_as_ubyte
from matplotlib.patches import Wedge
import matplotlib.colors as mcolors
image = data.astronaut()
image = Image.fromarray(img_as_ubyte(image))
# Function to allow performing the task
def adjust_hue(image, hue):
img_hsv = image.convert(’HSV’)
np_img = np.array(img_hsv)
hue_shift = int(hue * 255)
np_img = np_img.astype(np.int32)
np_img[..., 0] = (np_img[..., 0] + hue_shift) % 256
np_img = np.clip(np_img, 0, 255).astype(np.uint8)
adjusted_img = Image.fromarray(np_img, mode=’HSV’).convert(’RGB’)
return adjusted_img
# Function to create widgets
def create_hue_widgets():
# Slider
slider_label = widgets.Label(value=’Slider:’)
slider = widgets.FloatSlider(value=0.0, min=0.0, max=1.0, step=0.01)
# Dropdown
dropdown_label = widgets.Label(value=’Dropdown:’)
dropdown = widgets.Dropdown(options=[0.0, 0.2, 0.4, 0.6, 0.8], value=0.0)
# Radio Buttons
radio_buttons_label = widgets.Label(value=’Radio Buttons:’)
radio_buttons = widgets.RadioButtons(options=[0.0, 0.2, 0.4, 0.6, 0.8], value=0.0)
# Text Field
text_field_label = widgets.Label(value=’Text Field:’)
text_field = widgets.BoundedFloatText(value=0.0, min=0.0, max=1.0, step=0.01)
# Preset Buttons
preset_label = widgets.Label(value=’Preset buttons:’)
preset_hues = [
(0.0, ’red’),
(0.2, ’green’),
(0.4, ’cyan’),
(0.6, ’blue’),
(0.8, ’magenta’)
hue_mapping = {f"{hue}": hue for hue, color in preset_hues}
preset_buttons = [
widgets.Button(
description=f"{hue}",
layout=widgets.Layout(width=’75px’, height=’30px’),
style={’button_color’: color}
for hue, color in preset_hues
preset_buttons_box = widgets.HBox(preset_buttons)
# Color Wheel
color_wheel_label = widgets.Label(value=’Color Wheel:’)
color_wheel = widgets.Output()
create_color_wheel(color_wheel)
# Color Picker
color_picker_label = widgets.Label(value=’Color Picker:’)
color_picker = widgets.ColorPicker(concise=True, value=’#ffffff’, disabled=False)
# Layout
layout = widgets.Layout(align_items=’flex-start’)
spacer = widgets.Box(value=’’, layout=widgets.Layout(height=’20px’))
# Combine widgets into a vertical box
widgets_box = widgets.VBox([slider_label, slider, spacer,
dropdown_label, dropdown, spacer,
radio_buttons_label, radio_buttons, spacer,
text_field_label, text_field, spacer,
preset_label, preset_buttons_box, spacer,
color_wheel_label, color_wheel, spacer,
color_picker_label, color_picker, spacer], layout=layout)
return widgets_box, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, color_wheel, color_picker
# Function to create and display the color wheel
def create_color_wheel(output):
with output:
clear_output(wait=True)
fig, ax = plt.subplots(figsize=(1.5, 1.5))
num_colors = 360
theta = np.linspace(0, 2 * np.pi, num_colors, endpoint=False)
colors = plt.cm.hsv(theta / (2 * np.pi))
for i in range(num_colors):
wedge = Wedge(center=(0, 0), r=1, theta1=(i * 360 / num_colors),
theta2=((i + 1) * 360 / num_colors), color=colors[i],
transform=ax.transData._b, clip_on=False)
ax.add_patch(wedge)
ax.set_aspect(’equal’)
ax.set_xlim(-1.1, 1.1)
ax.set_ylim(-1.1, 1.1)
ax.axis(’off’)
fig.canvas.mpl_connect(’button_press_event’, on_color_wheel_click)
plt.show()
# Function to handle color wheel click
def on_color_wheel_click(event):
if event.inaxes:
x, y = event.xdata, event.ydata
theta = np.arctan2(y, x) % (2 * np.pi)
hue = theta / (2 * np.pi)
hue = round(hue, 2)
update_plot(None, image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, hue=hue)
# Function to convert hex color to hue value
def hex_to_hue(hex_color):
rgb = np.array([int(hex_color[i:i+2], 16) for i in (1, 3, 5)]) / 255.0
hsv = mcolors.rgb_to_hsv(rgb.reshape(1, 1, 3))
hue = hsv[0, 0, 0]
return hue
# Function to update the image display based on widget values
def update_plot(change, image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, hue=None):
with output:
clear_output(wait=True)
if hue is None:
if change and change[’owner’] in preset_buttons:
clicked_button = change[’owner’]
hue = hue_mapping[clicked_button.description]
elif change and change[’owner’] == slider:
hue = slider.value
elif change and change[’owner’] == dropdown:
hue = dropdown.value
elif change and change[’owner’] == radio_buttons:
hue = radio_buttons.value
elif change and change[’owner’] == text_field:
hue = text_field.value
elif change and change[’owner’] == color_picker:
hue = hex_to_hue(change[’new’])
else:
hue = 0.0
adjusted_image = adjust_hue(image, hue)
plt.figure(figsize=(4, 4))
plt.imshow(adjusted_image)
plt.axis(’off’)
plt.title(f’Hue Adjustment: {hue:.2f}’)
plt.show()
# Function to link widgets to the update function
def link_widgets_to_update(image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, color_picker):
def callback(change):
update_plot(change, image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping)
slider.observe(callback, names=’value’)
dropdown.observe(callback, names=’value’)
radio_buttons.observe(callback, names=’value’)
text_field.observe(callback, names=’value’)
color_picker.observe(callback, names=’value’)
for btn in preset_buttons:
btn.on_click(lambda btn: update_plot({’owner’: btn}, image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping))
# Function to display the widgets and output plot
def display_hue_widgets():
global output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, color_picker
output = widgets.Output()
widgets_box, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, color_wheel, color_picker = create_hue_widgets()
layout = widgets.Layout(width=’100%’, display=’flex’, justify_content=’space-between’)
ui_and_image_box = widgets.HBox([widgets_box, output], layout=layout)
display(ui_and_image_box)
link_widgets_to_update(image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping, color_picker)
update_plot(None, image, output, slider, dropdown, radio_buttons, text_field, preset_buttons, hue_mapping)
display_hue_widgets()
Appendix C User Evaluation Briefing and Instructions
Welcome!
This study evaluates user preferences for different UI widgets.
Your participation is voluntary. You are free to quit the study at any time. You can participate in this study only once.
All your responses (including demographic info) will be kept private and only used for research purposes.
# Study Instructions
- You will be working on 3 tasks. Each task is followed by 6 questions asking your UI widget preference by selecting one of two provided options for each question.
- For each task, please follow the task description and interact with **all the provided UI widgets**.
- After interacting with the UI widgets, you will choose the preferences that make the most sense to you regarding the widgets **predictability**, **efficiency**, or **explorability**.
- Predictability: Allows you to obtain results with no surprises or doesnt require you to deduce how to perform the interaction.
- Efficiency: Allows you to perform tasks with a minimum amount of effort.
- Explorability: Allows you to explore multiple possibilities and perform functions with high flexibility.
- When making your selection, please consider the following **comprehensive factors**:
- **Widget types**: Whether the widget types make sense to you regarding predictability, efficiency, or explorability.
- **Scores**: How much each widget is recommended in terms of predictability, efficiency, or explorability.
- Each widget type has a score out of 10.
- In each option, a widget with higher scores indicates a stronger recommendation.
- The total score of all widgets in each option adds up to 10.
- **Reasons**: Why each widget is a good choice to allow predictability, efficiency, or explorability.
- Some reasons may refer to users UI widget preferences when performing other relevant tasks. These reasons are based on a library. Assume that such a library exists.
- If a reason is based on the library, you will need to assess how well the information from this library is used for reasoning.
- **Do NOT** base your preference selection on:
- The number of widget types, e.g., choosing an option simply because it offers more widget types.
- The length of reasoning, e.g., choosing an option simply because the reasons are longer.
- Remember, your preference selection should consider widget types, scores, and reasons **collectively**. You can go back and interact with the UI widgets if you want.
- If widget types are the same, consider the scores and reasons to make your selection.
- If widget types and scores are the same, consider the reasons to make your selection.
- All questions are required, including preference selection and an explanation of your reason.

### v2

@@
-CrowdGenUI: Enhancing LLM-Based UI Widget Generation with a Crowdsourced Preference Library
+CrowdGenUI: Aligning LLM-Based UI Generation with Crowdsourced User Preferences
 Abstract.
-Large Language Models (LLMs) have demonstrated remarkable skills across various design domains, including UI generation.
-However, current LLMs for UI generation tend to offer generic solutions that lack a deep understanding of task context and user preferences in specific scenarios.
-We present CrowdGenUI, a framework that enhances LLM-driven UI generation with a crowdsourced user preference library.
-This approach addresses the limitations of existing methods by guiding LLM reasoning with user preferences, enabling the generation of UI widgets that align more closely with user needs and task-specific requirements.
-Using image editing as a test domain, we built this library from 50 users, capturing 720 user preferences, which include the predictability, efficiency, and explorability of multiple UI widgets.
-In a user study with 72 additional participants, our framework outperformed standard LLM-generated widgets in meeting user preferences and task requirements.
-We discuss these findings to inform future opportunities for designing user-centered and customizable UIs by comprehensively analyzing the extendability of the proposed framework and crowdsourced library.
+Large Language Models (LLMs) have demonstrated remarkable potential across various design domains, including user interface (UI) generation.
+However, current LLMs for UI generation tend to offer generic solutions that lack a nuanced understanding of task context and user preferences.
+We present CrowdGenUI, a framework that enhances LLM-based UI generation with crowdsourced user preferences.
+This framework addresses the limitations by guiding LLM reasoning with real user preferences, enabling the generation of UI widgets that reflect user needs and task-specific requirements.
+We evaluate our framework in the image editing domain by collecting a library of 720 user preferences from 50 participants, covering preferences such as predictability, efficiency, and explorability of various UI widgets.
+A user study (N=78) demonstrates that UIs generated with our preference-guided framework can better match user intentions compared to those generated by LLMs alone, highlighting the effectiveness of our proposed framework.
+We further discuss the study findings and present insights for future research on LLM-based user-centered UI generation.
 user preference, UI generation, large language models
-$c$s: Human-centered computing Collaborative and social computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing User studiesccs: Information systems Crowdsourcing
-Figure 1.
-CrowdGenUI is a framework that allows the generation of task requirement-aware and user preference-aligned UI widgets powered by LLMs.
-Based on input information specifying the content to edit, the task to perform, and various user preferences, the LLM-driven UI widget generation is enhanced by our crowdsourced user preference library together with a set of widget candidates for UI widget reasoning and code generation.
+ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing User studiesccs: Information systems Crowdsourcing
 1. Introduction
-Graphical user interfaces (GUIs) play a fundamental role in how most people operate computers and other computing devices.
-Carefully designed by expert interface designers, these GUIs aim to provide users with learnable ways to interact with software (Adobe, 2024c, b).
-Although effective, such interfaces can be complicated and sometimes overwhelming for users.
-In recent years, the emergence of large language models (LLMs) has led to increased interest in natural language interfaces.
-These language-based interfaces process human language input, analyze high-level intentions, and turn the interpreted intents into operations, providing a simplified way for users to interact with technology (OpenAI, 2024a; Descript, 2024; Maddigan and Susnjak, 2023).
-However, while language interfaces are promising, they are typically suited for high-level task execution, such as removing an object from an image or adding a filtering effect to videos.
-They often lack the precision required when users need to fine-tune specific parameters, such as adjusting the color in an image or customizing detailed settings in a video editing project.
-To address the complexity of professional GUIs and the inaccuracy of natural language interfaces, researchers have begun to explore the integration of LLMs with UI generation to combine the high-level task understanding enabled by LLMs with the precise, granular control offered by GUIs (Vaithilingam et al., 2024; Cheng et al., 2024).
-By turning natural language inputs into interactable UI widgets through UI generation, these systems provide an exciting hybrid interaction model that supports both user interaction with widgets for precise control and language as input for high-level task interpretation.
-Nevertheless, the existing UI generation systems also come with challenges.
-These LLM-generated UIs tend to be generic and may not fully account for the nuances of domain-specific tasks or the diverse preferences of individual users.
-Without specialized knowledge of interface design or a deeper understanding of user preferences, these UIs may fail to provide the optimal interaction experience.
-To overcome these challenges, we propose a novel framework, CrowdGenUI, that enhances LLM-driven UI generation by incorporating a crowdsourced user preference library, as illustrated in Figure 1.
-This framework leverages the collective insights of users to guide the LLM in generating tailored UI widgets for different user preferences.
-While the framework is broadly applicable across various domains, we demonstrate its effectiveness in the context of image editing within a Python programming environment.
-To build the preference library, we collected data from 50 users, with no requirement on UI design or programming experience, capturing their preferences across three key dimensions of interface design, i.e., predictability, efficiency, and explorability.
-By integrating this library into the LLM’s UI widget generation process, our system generates widgets that are not only more precise but also better aligned with user preferences and task-specific requirements.
-In our user study with a different set of 72 participants, we found that UI widgets generated based on our proposed framework were more aligned with user preferences and task requirements than those generic widgets generated by the LLM alone.
-This demonstrates the potential of our framework to improve LLM-based UI generation and suggests that crowdsourced user preference data can play a crucial role in creating user-centered user interfaces.
-To summarize, this paper makes the following key contributions:
-  • We propose a general framework to enhance LLM-driven UI generation by incorporating a crowdsourced user preference library for task-aware and preference-aligned widget generation.
-  • As a concrete implementation of this framework, we develop a system for the domain of image editing within a Python programming environment.
-This system includes crowdsourced data from 50 users, capturing user preferences for UI widget configurations in image editing tasks.
-  • We conduct a user study with an additional 72 participants to evaluate the effectiveness of the implemented system.
-The results of this study are extensively discussed to inform future research on the development of customizable and user-centered UIs.
+Graphical user interfaces (GUIs) have long served as the primary medium through which users interact with software applications and digital systems (Adobe, 2024c, b).
+Despite their ubiquity and effectiveness, designing intuitive and efficient GUIs remains a complex and labor-intensive process.
+It often requires expert knowledge, iterative user testing, and significant time investments—both from designers to refine the UI and from users to learn how to use it effectively.
+In response, there has been growing interest in natural language interfaces, which offer an alternative mode of interaction by allowing users to express their intentions through language.
+These interfaces reduce the design and learning overhead, simplifying how users interact with complex tools (OpenAI, 2024a; Descript, 2024; Maddigan and Susnjak, 2023).
+However, while convenient, natural language interfaces often fall short in scenarios requiring precision and control, such as adjusting image brightness or fine-tuning the parameters of a filter.
+Moreover, they typically lack the interactive, real-time feedback that GUIs provide, which is essential for tasks that benefit from trial-and-error and visual confirmation.
+With recent advances in generative AI, particularly the emergence of large language models (LLMs), researchers have started exploring how LLMs might bridge this gap by automatically generating adaptive user interfaces (Vaithilingam et al., 2024; Cheng et al., 2024).
+These systems aim to combine the high-level task understanding and domain flexibility of LLMs with the precise, user-driven control of traditional GUIs.
+Yet, current LLM-based UI generation methods rely mainly on the model’s internal reasoning and general knowledge.
+As a result, the generated UIs often appear generic and may fail to capture the nuanced requirements of specific tasks or the varied preferences of individual users.
+To address these limitations, we present CrowdGenUI, a novel framework that enhances LLM-based UI generation by incorporating crowdsourced user preferences.
+This framework leverages collective insights from real users to guide the LLM in generating UI widgets that are better tailored to different user preferences and task contexts.
+Although broadly applicable, we focus on the image editing domain within a Python-based environment to demonstrate the framework’s effectiveness.
+We constructed a preference library by collecting data from 50 participants to capture general users’ preferences across three core dimensions of UI design: predictability, efficiency, and explorability.
+This preference data is then integrated into the LLM’s UI generation process to augment LLM-based widget reasoning and code generation to produce UI widgets that align with user and task demands.
+In a user study with 78 additional participants, we found that UI widgets generated using our framework were rated as more aligned with user preferences and task requirements compared to those generated by the LLM alone.
+These results highlight the potential of our framework to advance user-centered interface generation and demonstrate that integrating crowdsourced preferences can significantly enhance the effectiveness of LLM-based UI generation.
+In summary, our main contributions are as follows:
+  • We present a framework that enhances LLM-based UI generation by incorporating crowdsourced user preferences to support task-aware and preference-aligned UI widget generation.
+We introduce the framework design and the methodology for crowdsourcing user preference data.
+  • We implement a system prototype for this framework in the image editing domain.
+This implementation delivers 720 user preferences crowdsourced from 50 users and a structured reasoning technique to guide LLM reasoning and code generation for user preference-aligned UI generation.
+  • We conduct a user study with an additional 78 participants that discovers the impact of crowdsourced user preference data on UI generation.
+By discussing the study findings, we provide insights for future research on creating adaptive and user-centered LLM-generated UIs with user data as guidance.
 2. Related Work
-To provide the background for our work, we introduce prior work on traditional fixed UI, natural language UI, and dynamically generated UI.
-We discuss their advantages and limitations and compare them with our proposed framework.
-A visual comparison is presented in Figure 2.
-2.1. Traditional Fixed UI
-Traditional UI design in professional digital software like Illustrator (Adobe, 2024b), Photoshop (Adobe, 2024c), Premiere Pro (Adobe, 2024e), and Blender (Blender, 2024) follows a designer-centric model, where a fixed set of tools and controls are designed for various tasks.
+To provide background for our work, we introduce prior research on traditional fixed UIs, natural language UIs, and dynamically generated UIs, and discuss their advantages as well as limitations.
+2.1. Traditional Fixed UIs
+Traditional UI design in professional digital software like Illustrator (Adobe, 2024b), Photoshop (Adobe, 2024c), Premiere Pro (Adobe, 2024e), and Blender (Blender, 2024) follows a designer-centric model, where a fixed set of tools and UI controls are designed for various tasks.
 These applications rely on well-established interface elements such as dropdown menus, sliders, text fields, and panels that allow users to manipulate images, graphics, videos, and 3D models.
@@
 Premiere Pro delivers video editing features through timelines and control panels, and Blender offers a complex interface for 3D modeling and animation.
-These tools are powerful but often rigid, requiring users to adapt to the software’s structure.
-As a result, their complexity can be overwhelming to users, and their interface may not always align with individual workflows or preferences, which limits the flexibility and personalization of user experience.
-$2.$ Natural Language UI
+These tools are powerful yet rigid, requiring users to learn and adapt to the software’s structure.
+As a result, their complexity can overwhelm beginner users, and their interface, though it has undergone interactive user testing, may not always align with individual workflows or preferences, limiting the flexibility and personalization of the user experience.
+2.2. Natural Language UIs
 Using natural language as input to edit digital content has gained increasing interest due to its simplicity.
-This approach allows users to perform tasks like image, video, data visualization, 3D models, and human motion editing by simply describing the changes they want rather than navigating complex software interfaces.
-For example, natural language-based image editing tools, such as DALL-E (OpenAI, 2024a) or Runway (Runaway, 2024), allow users to generate and edit images by typing commands like “make the sky bluer” or “add a tree in the background”.
-Similarly, tools like Adobe Firefly (Adobe, 2024a) and Photoshop Generative Fill (Adobe, 2024d) enable image generation and edits using prompts like “change to a vintage look” or “add a yellow bee in the circled region next to the flower”.
-In video editing, tools like Descript (Descript, 2024) allow users to edit video content by editing a transcript, and tools like Runway and Premiere Pro with Sensei AI (Adobe, 2024f) also support video editing using natural language prompts like “add a motion blur effect”.
-Also, there are many research projects (Brooks et al., 2023; Zhang et al., 2024; Fu et al., 2024, 2022; Qin et al., 2024) that have significantly advanced the techniques for image and video editing using language as the primary modality.
-In the field of visualization, systems like Chat2VIS (Maddigan and Susnjak, 2023), NL4DV (Sah et al., 2024) and Data Formulator (Wang et al., 2023, 2024) allow users to create charts and graphs by simply asking for a “bar chart showing sales over time”, and LIDA (Dibia, 2023) enable the generation of grammar-agnostic visualizations and infographics using LLMs.
-For the generation and editing of 3D model (Chang et al., 2017; Poole et al., 2022) and human motion (Tevet et al., 2022; Raab et al., 2023; Chen et al., 2023; Liu and Sra, 2024), prior research has investigated using language to describe users’ goals, such as “create a 3D lemur taking notes in a journal” or “change the upper body motion to throwing a ball and keep the original running motion”.
-These innovations are transforming traditional workflows by making advanced editing more accessible to non-expert users, reducing the technical execution gulf and enabling them to focus on creative intent.
-However, natural language interfaces struggle with precise control over specific details and cannot effectively support interactive editing, limiting their ability to handle complex, fine-grained adjustments typically required in professional workflows.
-2.3. Dynamically Generated UI
+This approach allows users to perform tasks like the generation and editing of images (OpenAI, 2024a; Runaway, 2024; Adobe, 2024a, d; Brooks et al., 2023), videos (Descript, 2024; Adobe, 2024f; Qin et al., 2024), data visualizations (Maddigan and Susnjak, 2023; Sah et al., 2024; Wang et al., 2023, 2024; Dibia, 2023), 3D models (Chang et al., 2017; Poole et al., 2022), and human motions (Tevet et al., 2022; Raab et al., 2023; Chen et al., 2023; Liu and Sra, 2024b, a) by simply describing the outputs they want rather than navigating complex software interfaces.
+Natural language UIs are transforming traditional workflows by making advanced editing more accessible to non-expert users, reducing the technical execution gulf, and enabling users to focus on creative intent.
+However, these UIs require articulation and can hurt usability as half the population is insufficiently articulate in writing to use natural language UIs well (Nielsen, 2023).
+Natural language UIs also struggle with precise control over specific details.
+They cannot effectively support interactive editing, which limits their ability to handle complex, fine-grained adjustments and provide prompt visual feedback of intermediate results typically required in iterative workflows.
+2.3. Dynamically Generated UIs
 Recent research has studied generating dynamic UIs to provide users with precise, direct interactions for content editing.
 One prominent approach is the relaxation method, which involves creating UI widgets that provide generalized control over variables within a function or query.
-For instance, Mavo (Verou et al., 2016) provides users with the capability to construct interactive HTML pages by incorporating unique attributes, which, in turn, recommend editing widgets tailored to those attributes.
+For instance, Mavo (Verou et al., 2016) allows users to construct interactive HTML pages by incorporating unique attributes, which, in turn, recommend editing widgets tailored to those attributes.
 Heer et al.
-(Heer et al., 2008) employ query relaxation techniques, allowing users to broaden their selection effectively.
+(Heer et al., 2008) employ query relaxation techniques, allowing users to broaden their selection tasks effectively.
 Bespoke (Vaithilingam and Guo, 2019) utilizes user demonstrations to generate custom UIs for command-line applications.
 It applies rule-based heuristics to infer the semantic types of parameters in bash commands and thus dynamically create widgets for parameter manipulation.
-Another line of research, known as precision interfaces, uses SQL queries as a framework for UI generation (Zhang et al., 2018; Chen, 2020; Chen and Wu, 2022).
+Another line of research, known as precision interfaces, uses SQL queries as a framework for UI generation (Zhang et al., 2018; Chen, 2020; Chen and Wu, 2022; Chen et al., 2022).
 These systems convert a sequence of input queries into interactive widgets, allowing users to adjust query parameters directly.
-An example is NL2Interface (Chen et al., 2022), which translates natural language commands into SQL queries and subsequently generates a UI for users to edit parameters within the queries.
-In addition to generating dynamic UIs from command-line applications and queries, recent advancements have focused on creating UIs directly from natural language descriptions.
-This approach allows users to articulate their goals in natural language, and the system generates an interactive UI tailored to the task at hand using LLM’s reasoning capabilities.
+In addition to generating UIs from command-line applications and SQL queries, recent advancements have focused on creating UIs from natural language.
+This approach allows simplicity by taking users’ natural language as input, and the system generates an interactive UI tailored to users’ task descriptions using LLM’s reasoning and code generation capabilities.
 For instance, DynaVis (Vaithilingam et al., 2024) generates dynamic interfaces for visualization editing based on natural language commands.
-By enabling users to describe changes to their visualizations through simple text, such as “rotate the x-labels by 30 degrees counter-clockwise”, DynaVis dynamically creates suitable widgets, such as a slider for adjusting the rotation of x-labels.
+By enabling users to describe changes to their visualizations through simple text, such as “rotate the x-labels by 30 degrees counter-clockwise”, DynaVis dynamically creates suitable widgets, such as a slider, for adjusting the rotation of x-labels.
 Similarly, Biscuit (Cheng et al., 2024) generates UIs to help users follow machine learning tutorials in Jupyter Notebooks using natural language queries.
-Users can describe tasks like “sample training data” or “adjust model hyperparameters”, and Biscuit generates ephemeral UIs for the specified parameter adjustments.
-Although natural language-based UI generation generates UIs without depending on fixed rules or query relaxation techniques and utilizes LLMs’ advantage to interpret intents from high-level descriptions, most prior work using standalone LLMs does not fully consider individual user preferences and domain-specific task requirements.
-This can result in interfaces that, while functional, may not be the most effective or usable for specific users, especially when personalization and task-specific requirements are needed.
-3. User Feedback Collection and Principles for UI Design
-While LLM-powered UI generation offers many advantages, it primarily relies on generic training data, lacking task-specific contexts and user-centered insights, such as individual preferences.
+Users can describe tasks like “sample training data” or “adjust model hyperparameters”, and Biscuit generates UIs for the specified parameter adjustments.
+Although natural language-based UI generation supports both precision with interactable UI widgets and simplicity by taking natural language input, most prior work using standalone LLMs does not fully consider individual user preferences and domain-specific task requirements.
+This can result in generated interfaces that, while functional, may not be the most effective for specific users, especially when personalization and task-specific requirements are needed.
+Figure 1.
+Overview of the CrowdGenUI framework.
+This framework consists of four main components: it takes (1) user context as input, including task description and UI control preferences, and this information is passed to the LLM together with (2) a user preference library and (3) a collection of UI widget candidates.
+All the input data is used to guide the LLM reasoning process to obtain (4) generated UI widgets that are aligned with user preferences and task requirements.
+[…]
+This figure is a flow chart that takes user context, user preference library, and UI widget candidates as input for LLMs, and generates UI as the output.
+3. CrowdGenUI: A Framework that Aligns UI Generation with User Preference
+3.1. Overview
+Figure 1 provides an overview of our proposed framework, CrowdGenUI, which aligns generated UI widgets with task requirements and specific user preferences by incorporating a user preference library that captures users’ preferred UI widgets and their rationale to guide LLM-based UI generation.
+Enhanced alignment through user preference integration: Unlike prior work on LLM-based UI generation, this framework integrates user preference insights directly into the LLM reasoning process.
+The user insights consist of crowdsourced preference data for UI widgets from real users.
+By designing a systematic reasoning technique to guide LLM reasoning, this framework is designed to boost LLM alignment in the UI generation application.
+Flexibility and extendability through modular design: By using a modular-based architecture, this framework allows independent and flexible module updates to accommodate extendable contexts.
+This applies to all three input modules: (1) user context can evolve with updated task requirements and various preferences, ensuring continual alignment with specific user needs; (2) user preference library can continuously take updated user data, expanding the diversity and richness of user insights used during UI generation; (3) UI widget candidates can flexibly incorporate either static widget sets from established libraries or dynamically produced widgets, thereby expanding potential UI outcomes.
+3.2. Usage Scenario
+Before diving into the technical implementation details, we guide readers through an example usage scenario, shown in Figure 2, to demonstrate how CrowdGenUI can be used for a content editing task.
+Imagine that Alice wants to change the tone of an image from green, spring colors to warm, fall colors.
+She uses CrowdGenUI to edit her image through the following process:
+Figure 2.
+Example usage scenario of CrowdGenUI.
+A user wants to change the color of an image to fall colors.
+They provide the image, describe the task, and specify their UI control preference.
+Taking user input, CrowdGenUI reasons the proper UI widgets and generates the code for widget implementation.
+The user uses the generated widgets to edit the image and can further update their input to obtain updated UIs, e.g., from efficient experiments using preset options to fine-grained control using a slider.
+[…]
+This figure demonstrates an example usage scenario using a flow chart. From left to right, the chart consists of interactive user input, UI widget reasoning and implementation, and content editing.
+Step 1—Describe task and specify preferences: She describes her needs as “Change this image to fall colors” and provides the image she wants to edit.
+She also specifies that she prefers to experiment with image colors quickly with color tone overviews.
+Step 2—Use CrowdGenUI for UI widget reasoning and implementation: CrowdGenUI processes Alice’s input and determines the most suitable UI widgets for her task.
+Since the task involves adjusting the image’s color tone efficiently, CrowdGenUI consults the crowdsourced user preference library to identify relevant tasks and commonly preferred UI widgets for efficiency.
+Based on this information, it reasons that preset buttons displaying color tone previews are the optimal choice.
+Next, it generates code to implement an image tone adjustment function and a preset button function, links the preset buttons to image tone adjustment, and returns the code for Alice to execute and obtain the widgets.
+Step 3—Edit content with generated UI: Alice interacts with the preset buttons to explore various pre-defined tones.
+The color previews on buttons help her to anticipate how each selection will affect the image, and the preset values enable quick experimentation with simple button clicks.
+Step 4—Update requirements for iterative UI generation: Although the preset buttons satisfy Alice’s initial needs, she finds an orange tone she likes but wants to make it slightly redder.
+Since the preset options do not allow that, she updates her request by asking for UI widgets that support finer tone adjustments.
+This brings her back to Step 2, where CrowdGenUI generates new widgets based on her updated requirements, and moves on to Step 3, where she continues refining her image using the newly generated widgets.
+4. CrowdGenUI Implementation
+Figure 3 illustrates a system prototype we implemented for our proposed framework. In this section, we dive into the technical details and implementation rationale.
+Figure 3.
+System implementation pipeline.
+The system starts with natural language input specifying the task description and user preferences.
+With this input, the LLM applies chain-of-thought (CoT) reasoning to obtain reasoned widgets.
+Based on this reasoning process, the crowdsourced user preference library, containing the crowdsourcing tasks, user-preferred widgets, and their rationale, is integrated to augment widget reasoning.
+Additionally, the Jupyter Widget Library widget candidates are fed to the LLM for reasoned widget implementation.
+Lastly, the reasoned widgets and reference code snippets are passed to the LLM for code generation in Python.
+This allows widget implementation and interaction in Jupyter Notebook.
+[…]
+This figure is a flow chart showing the system implementation pipeline.
+Starting from the left, the task description, preference aspect, crowdsourced preference library, and UI widget candidates are passed as input to the LLM for CoT reasoning.
+As the reasoning output, reasoned widgets containing widget types and the corresponding rationale are returned.
+This output is further passed to the LLM for code generation with sample code as guidance.
+Finally, Python code is returned for widget display and interaction.
+4.1. User Context
+The system input is user context, which consists of task description and UI control preferences.
+4.1.1. Task Description
+The task description needs to specify the content to edit (e.g., an image), values to adjust (e.g., image color), and goals to achieve (e.g., changing the image to fall colors).
+4.1.2. UI Control Preferences
+UI control preferences can involve various preference aspects. In the current system, we use predictability, efficiency, and explorability to implement our prototype.
+These preferences are chosen based on well-established UI design principles developed by prior research.
+Specifically, McKay’s UI is Communication (McKay, 2013) emphasizes that an intuitive interface relies on effective communication that balances multiple principles, including predictability, efficiency, discoverability, and explorability.
+McKay argues that intuitive UIs align with a user’s prior experiences, providing clear affordances and responsive feedback to guide interaction.
+Lidwell et al.’s Universal Principles of Design (Lidwell et al., 2010) expands on these ideas by highlighting fundamental design principles such as the flexibility-usability tradeoff, mapping, and mental models, which underscore the importance of aligning system controls with user expectations and simplifying complex interactions.
+Together, these principles highlight that successful UIs are predictable and efficient and allow for exploratory interaction without overwhelming the user, leading to our selection of predictability, efficiency, and explorability to guide LLM-based UI generation.
+We introduce the user preference library that we crowdsourced according to the three aspects in Section 5.
+4.1.3. Input to the LLM
+Apart from feeding the image to the LLM, the rest of the user context is represented as text input denoted by task-description ($T$) and preference-aspect ($A$) in JSON format.
+4.2. UI Widget Reasoning
+User context is passed to the LLM for widget reasoning.
+Additional input for LLM reasoning includes the crowdsourced user preference library and UI widget candidates.
+Crowdsourced-library ($L$) and widget-candidates ($C$) are also text data formatted as JSON.
+With all the input data, we instruct the LLM to use a chain-of-thought (CoT) reasoning technique to reason UI widgets.
+We present the CoT reasoning overview in Algorithm 1 and the detailed prompt in Appendix B.1.
+The LLM first analyzes the user’s task description to identify the task requirements.
+Based on this analysis, it is instructed to analyze the similarity between the given task and the crowdsourcing tasks by comparing task descriptions to extract similar tasks.
+Once relevant tasks are identified, the LLM is further prompted to look for user-preferred widgets for each relevant task according to the preference aspect and the reasons users provided for their choices.
+It needs to refer to the located crowdsourced data to reason the most appropriate widgets that belong to the widget candidates.
+Finally, the UI widget reasoning results are formatted as JSON, including the user task’s name, the reasoned widgets for each preference aspect, and the corresponding reasoning.
+Algorithm 1 Overview of CoT Reasoning
+  Input: task-description ($T$), preference-aspect ($A$), crowdsourced-library ($L$), widget-candidates ($C$)
+  Output: widget-reasoning
+1:  for each crowdsourced-task in $L$ do
+2:     relevant-tasks $\leftarrow$ similarity ($T$, crowdsourced-task) {Identify relevant tasks}
+3:     for each aspect in $A$ do
+[…]
+5:        return  ui-widgets $\in$ $C$Return matching widgets}
+6:     end for
+7:  end for
+[…]
+4.3. UI Widget Implementation
+Once the reasoned widgets are obtained, the LLM is further used to generate the Python code to implement these widgets.
+During this process, we provide an example code, which covers the implementation of all the widget candidates, as guidance for the LLM to reference during code generation.
+Our experiment shows that including example code significantly reduces bugs in the LLM-generated code.
+More importantly, the example code enables the LLM to successfully generate widget implementations that go beyond the default functionality of the Jupyter Widgets Library.
+For instance, a color wheel widget requires additional use of the Interactive Canvas in Jupyter (Jupyter, 2024a) and the implementation of mouse interactions on the wheel to map clicks to the corresponding colors.
+In addition to specific widget implementations, the example code also offers a structure to organize and display widgets in a clear, structured format within Jupyter Notebook for user interaction.
+The detailed prompt and example code for UI widget implementation can be found in Appendix B.2.
+4.4. System Implementation Rationale
+Task domain: We implement the system in the image editing domain because it offers generalizability across a diverse range of content editing domains.
+We carefully design image editing tasks that involve various data manipulation needs, including continuous and discrete value adjustments and tasks emphasizing color and position visual feedback (Section 5.3.2).
+These tasks require both precise control and iterative exploration, which are essential for content editing requirements in general.
+Moreover, the interactions employed in image editing, characterized by real-time visual feedback, trial and error, and fine-tuning content attributes, are directly transferable to other digital content editing applications such as video editing, data visualization, and graphic design.
+Consequently, image editing tasks address fundamental data manipulation needs that apply to broader digital content editing practices.
+User interface: The current implementation uses Jupyter Notebook to build the user interface as it can handle all the data manipulation requirements for our tasks.
+Although this implementation does not cover other UI possibilities, such as web or professional software, the framework is extendable to accommodate these UIs by updating the widget candidates to web or software widget libraries and generating code in the corresponding programming languages for widget implementation.
+UI widget candidates: The current implementation focuses on the Jupyter Widgets Library widgets rather than open-ended options.
+This allows practical feasibility since implementing an infinite variety of widgets can be unmanageable.
+Also, these widgets are strategically chosen as foundational elements that map to universal interaction principles.
+For instance, sliders and text fields for continuous adjustments can extend to more granular controls like range sliders or numeric steppers, and dropdowns and radio buttons for discrete selections can evolve into multi-select interfaces or dynamic filters.
+Focusing on these core widgets can allow scalability by integration with external libraries (e.g., Angular (Material, 2024), React Bootstrap (Bootstrap, 2024), or Vuetify (Vuetify, 2024)) to expand into advanced UI components in future iterations.
+LLM selection: We use GPT-4o (OpenAI, 2024b) for UI widget reasoning and implementation due to its advanced reasoning capabilities, robust code generation performance, and strong alignment with natural language instructions, which are critical for dynamically generating UIs based on task descriptions and user preferences.
+5. User Preference Crowdsourcing
+In this section, we introduce our collected user preference data to build the library for UI generation.
+We present why we use crowdsourcing to collect this data, followed by the crowdsourcing setup and an analysis of the collected data.
+Figure 4. Crowdsourcing user interface for task: adjusting an image to fall colors (image_adjust_fall_color).
+[…]
+This figure shows an example of the crowdsourcing user interface consisting of three components.
+From top to bottom, the components are the task description, the UI widget interaction panel to perform the task, and the user preference selection panel.
+The task description specifies the goal that the user is expected to achieve.
+Below it, the UI widget interaction panel shows multiple widgets on the left and an image to edit on the right.
+At the bottom, the preference selection panel allows users to select their preferred widgets and write their reasons for the three preference aspects.
+5.1. Motivation
 User preference is part of user feedback and plays a crucial role in UI design and UI widget selection.
 The user feedback data is traditionally collected through labor-intensive methods like interviews, usability tests, A/B tests, and feedback forms (Stone et al., 2005; Wilson, 2013).
-Given that LLMs are trained on large-scale human data, they provide a promising solution to incorporating user preferences into dynamically generated UIs.
-However, without dedicated user preference datasets, current LLMs cannot fully address this need.
-To fill this gap, we propose crowdsourcing a preference library to enhance LLM-driven UI widget generation with user preference data, motivated by the effectiveness of adopting crowdsourcing for user feedback collection to assist various design areas (Xu et al., 2014, 2015; Luther et al., 2015; Yu et al., 2016; Foong et al., 2017; Lee et al., 2018; Oppenlaender et al., 2020).
-To determine the specific aspects of user preference, we refer to well-established UI design principles developed by prior research.
-McKay’s UI is Communication (McKay, 2013) emphasizes that an intuitive interface relies on effective communication that balances multiple principles, including predictability, efficiency, discoverability, and explorability.
-McKay argues that intuitive UIs align with users’ prior experiences, providing clear affordances and responsive feedback to guide interaction.
-Lidwell et al.’s Universal Principles of Design (Lidwell et al., 2010) expand on these ideas by highlighting fundamental design principles such as the flexibility-usability tradeoff, mapping, and mental models, which underscore the importance of aligning system controls with user expectations and simplifying complex interactions.
-Together, these principles highlight that successful UIs are predictable and efficient and allow for exploratory interaction without overwhelming the user.
-Drawing from these foundational principles, we have identified predictability, efficiency, and explorability as three critical aspects to guide our approach to building a user preference library to enhance LLM-powered UI widget generation.
-The remainder of the paper is organized as follows: first, we provide an overview of our proposed framework.
-Next, we describe the implemented system in the image editing domain and discuss the crowdsourced data collection process.
-This is followed by a detailed user evaluation, along with a discussion of the limitations, potential future work, and the conclusion.
-Figure 2.
-Overview of our proposed CrowdGenUI framework and comparison with prior work.
-Our framework consists of four main components: it takes (1) task description in natural language as input, and this information is passed to the LLM together with (2) the collected user preference library and (3) a collection of UI widget candidates.
-All the input data is used to guide the LLM reasoning process to obtain (4) generated UI widgets that are aligned with user preferences and task requirements.
-4. Framework Overview
-Figure 2 provides an overview of our proposed framework, CrowdGenUI, and a visual comparison with previous research.
-Our framework improves existing UI widget generation systems that take task descriptions as input for LLMs to generate UIs dynamically.
-To align generated UI widgets with specific user needs, we collect a user preference library that captures users’ preferred UI widgets and their rationale.
-This library is provided to the LLM alongside task descriptions and widget candidates to guide the UI generation process.
-For implementation, the widget candidates are selected from the widget list of Jupyter Widgets Library (Jupyter, 2024b).
-These candidates can be easily extended to include additional widget options, such as those from other widget libraries or LLM-generated widgets.
-5. Crowdsourced Library-enhanced UI Widget Generation for Image Editing
-Figure 3 illustrates the system we implemented for our proposed framework.
-In this section, we dive into the technical details of this system, including UI widget reasoning, integration of the crowdsourced library into the reasoning process, and UI widget code generation.
-[…]
-5.1. UI Widget Reasoning
-[…]
-5.2. Integration of Crowdsourced User Preference Library
-[…]
-5.3. UI Widget Code Generation
-[…]
-5.4. Implementation Details
-We use GPT-4o (OpenAI, 2024b) as the LLM for both widget reasoning and code generation, with Python as the programming language to implement the widgets.
-Occasionally, the code generated by GPT-4o may have minor bugs and might not compile correctly.
-We manually fix these issues to ensure all code generated by the LLM runs properly during user evaluation.
-This setup aligns with our focus on LLM-driven widget reasoning and code generation rather than real-time UI generation.
-We use this system to serve as a prototype to demonstrate how LLM reasoning, when enhanced by a crowdsourced library, can effectively guide UI widget generation according to task requirements and user preferences, particularly in the image editing domain.
-6. Crowdsourcing a Preference Library for UI Widgets
-In this section, we introduce how we collect user preference data to build the library that enhances LLM-driven UI widget generation.
-Following our previous implementation, we gather preference data in the image editing domain to illustrate the process.
-We present the study setup and provide an analysis of the collected data.
-(a) Task description and provided UI widgets to perform the task.
-(b) UI widget preference selection panel.
-Figure 4. Crowdsourcing user interface for task: adjusting an image to fall colors (image_adjust_fall_color).
-6.1. Platform and Participants
-[…]
+Given that LLMs are trained on large-scale human data, they provide a promising solution to incorporating user feedback, such as user preferences, into dynamically generated UIs.
+However, without dedicated user preference datasets, current LLMs cannot fully address this need (recall Section 2.3).
+To fill this gap, we propose crowdsourcing a preference library to enhance LLM-based UI widget generation with user preference data, motivated by the effectiveness of adopting crowdsourcing for user feedback collection to assist various design areas in previous research endeavors (Xu et al., 2014; Hossain and Kauranen, 2015; Xu et al., 2015; Luther et al., 2015; Yu et al., 2016; Foong et al., 2017; Lee et al., 2018; Oppenlaender et al., 2020).
+5.2. Participants
+The crowdsourcing took place on Prolific (Prolific, 2024).
+We recruited participants with Python programming experience only to ensure they were familiar with running the provided Python code in Jupyter Notebook.
+During the study, they did not write code but just executed our provided code to interact with UI widgets and work on their tasks on desktops using a mouse and keyboard.
+We did not require participants to have image editing or UI programming experience, as we aim to collect regular users’ preference data and make generated UIs align with end users’ preferences.
+We acknowledge that this participant recruitment may limit our sample by excluding users who are unfamiliar with Python.
+However, these participants can still meaningfully contribute to general users’ UI preferences, as programming expertise is independent of UI usage and preference for non-programming tasks.
+According to prior research on user testing for UX design (Tomczak et al., 2023; University, 2025; Interviews, 2025), 20–30 participants are generally sufficient for evaluating simple to intermediate tasks, while 50 participants are recommended for more complex tasks.
+Following these established guidelines, we invited 50 participants to the crowdsourcing study (Age: 29.86 $\pm$ 7.83, Male: 37, Female: 12, Non-binary: 1), and compensated them 10 USD for the 30-minute study.
 5.3. Crowdsourcing Setup
 5.3.1. User Interface
-In the crowdsourcing study, we offered each participant a Jupyter Notebook to work on eight image editing tasks, interact with several UI widgets to perform each task and provide their preferences for the widgets.
-Figure 4 shows the user interface for a crowdsourcing task—adjusting an image to fall colors.
+During the crowdsourcing study, each participant had a Jupyter Notebook to work on eight image editing tasks, interact with multiple UI widgets to perform each task, and provide their preferences for the widgets.
+Figure 4 shows the UI for a crowdsourcing task—adjusting an image to fall colors.
 The interface contains the task requirement, multiple UI widgets, an image to edit, and a user preference selection panel.
@@
 The crowdsourcing tasks cover three categories in the image editing domain.
-We introduce these categories and provide an overview of the tasks below.
 Some tasks may overlap across multiple categories.
-Our goal is to use these tasks to incorporate a range of value adjustments and data manipulation needs in image editing.
-  • Continuous value adjustment: These tasks need to adjust a value across a continuous range smoothly. Examples include adjusting image lightness, saturation, hue, and color balance.
+Our goal is to use these tasks to incorporate a range of value adjustments and data manipulation needs.
+Continuous value adjustment: These tasks need to adjust a value across a continuous range. Examples include adjusting image lightness, saturation, hue, and color balance.
 Discrete value selection: These tasks involve selecting from predefined options. Examples include choosing predefined values for image lightness, saturation, hue, and watermark positions.
@@
 We use these tasks to encourage exploration, such as in Tasks 1, 2, 3, 6, and 7, while also emphasizing precision, as seen in Tasks 4, 5, and 8.
-Two tasks are used as control tasks: image_adjust_fall_color and image_color_match.
-We instruct participants to ensure that the final images best meet the task requirements and use these tasks as attention checks to help identify untrustworthy responses.
-We randomize the order of tasks, ensuring that no particular task appears consistently in the same position to reduce order effects, such as fatigue or priming, that could influence user preferences and behaviors.
+Tasks 4 and 5 are used as control tasks for attention checks to help us identify untrustworthy responses, where we instruct participants to ensure that the final images best meet the task requirements.
+We randomize the order of tasks, ensuring that no particular task appears consistently in the same position to reduce order effects that could influence user behaviors.
 Table 1. Crowdsourcing task names and descriptions.
-Task Name Task Description
-1 image_adjust_lightness Experiment with lightness settings to see how different levels affect the image.
+[…]
+This table consists of two columns: task name and task description.
+In each row, the number, name, and description of each crowdsourcing task are detailed.
+Task Name Task Description 1 image_adjust_lightness Experiment with lightness settings to see how different levels affect the image.
 2 image_adjust_saturation Boost the saturation to make the colors pop.
 3 image_adjust_hue Experiment with different hues to find a color tone that complements the overall mood of the image.
-4 image_adjust_fall_color Adjust the color settings to gradually shift the image tones, creating a warm, autumnal atmosphere. Ensure the final image is the one that looks the most autumnal to you.
-5 image_color_match Change the rocket’s color to match the provided reference color. Ensure the rocket’s color in the final image best matches the reference color.
+4 image_adjust_fall_color Adjust the color settings to gradually shift the image tones, creating a warm, autumnal atmosphere.
+Ensure the final image is the one that looks the most autumnal to you.
+5 image_color_match Change the rocket’s color to match the provided reference color.
+Ensure the rocket’s color in the final image best matches the reference color.
 6 image_adjust_color_balance Experiment with different color balance settings to see how altering the red, green, and blue levels affects the overall color harmony of the image.
@@
 5.3.3. User Preference Aspects
-We use three key aspects of UI widget preferences, drawing from prior research on intuitive user interface design principles (recall Section 3).
-These aspects—predictability, efficiency, and explorability—reflect diverse user preferences when interacting with digital software, with the detailed descriptions below:
-  • Predictability: Users can achieve results without encountering unexpected outcomes or needing to figure out how to perform the interaction.
-  • Efficiency: Users can perform tasks with minimal effort.
-  • Explorability: Users can explore multiple possibilities and perform tasks with high flexibility.
-6.2.4. UI Widget Selection and Implementation
-UI widget selection needs to consider task and data manipulation requirements, such as input and output data types, and match the widget types with the specified task requirements.
-We choose UI widgets for the crowdsourcing tasks from the Jupyter Widgets Library (Jupyter, 2024b), which offers both basic and advanced UI widgets, covering a wide range of control needs.
-Specifically, for continuous value adjustment tasks, sliders [Uncaptioned image] are typically ideal to allow for fine-grained control, and text fields [Uncaptioned image] can be used for precise value input.
-For discrete value selection tasks, dropdown menus [Uncaptioned image] and radio buttons [Uncaptioned image][Uncaptioned image] are often effective.
-For tasks emphasizing visual feedback, UI widgets like color pickers [Uncaptioned image], color wheels [Uncaptioned image], direct manipulation (e.g., clicking [Uncaptioned image]), and preset buttons with preview overlays [Uncaptioned image][Uncaptioned image] allow users to see outcome previews through color or position visual cues.
-In this work, we present one approach to UI widget implementation and selection, acknowledging that it does not cover all possible options.
-However, this method can be easily expanded by leveraging additional UI widget libraries, such as (Material, 2024; Bootstrap, 2024; Vuetify, 2024).
-These libraries can facilitate data manipulation and content editing in software or websites and enable the inclusion of a broader range of widget types in various applications.
-6.3. Crowdsourced UI Widget Preference Library
-6.3.1. Overview
-We distributed the crowdsourcing study to the 50 participants in three versions.
+We use predictability, efficiency, and explorability as user preference aspects, drawing from prior research on intuitive user interface design principles (recall Section 4.1.2).
+(1) Predictability emphasizes that users can achieve results without encountering unexpected outcomes or needing to figure out how to perform the interaction.
+(2) Efficiency highlights that users should be allowed to perform tasks with minimal effort.
+(3) Explorability underscores that users can explore multiple possibilities and perform tasks with high flexibility.
+5.3.4. UI Widget Candidates
+As presented in Section 4.4, we aim to select UI widgets to cover various data manipulation requirements and align these candidates with the specified task requirements.
+To this end, the Jupyter Widgets Library (Jupyter, 2024b) offers both basic and advanced UI widgets, including a wide range of control needs.
+Specifically, for continuous value adjustment tasks, sliders are selected to allow for fine-grained control, and text fields can be used for precise value input; for discrete value selection tasks, dropdown menus and radio buttons are often effective; for tasks emphasizing visual feedback, UI widgets like color pickers, direct manipulation (e.g., clicking), and preset buttons with preview overlays are picked to show outcome previews through color or position cues.
+5.4. Crowdsourced UI Preference Library
+5.4.1. Statistics Overview
+We distributed the crowdsourcing study to our participants in three versions.
 Each version included all eight tasks and a subset of the three preference aspects, covering one, two, or all three aspects.
-This approach ensured that each participant provided their preferences for 15-18 preference selections out of the total 3 aspects $\times$× 8 tasks = 24 selections across all tasks, helping to prevent fatigue.
+This approach ensured that each participant provided their preferences for 15-18 preference selections out of the total 3 aspects $\times$ 8 tasks = 24 selections across all tasks, helping to prevent fatigue.
 The study yielded a UI widget preference library with 30 unique user responses for each aspect.
-In total, we obtained 30 responses $\times$× 3 aspects $\times$× 8 tasks = 720 user preferences.
-Figures 5 and 6 show the distribution of user preferences for all crowdsourcing tasks.
-Figure 5. User preference distribution of tasks 1-4 collected in the crowdsourcing study.
-Figure 6. User preference distribution of tasks 5-8 collected in the crowdsourcing study.
-6.3.2. Analysis of User Preference Distribution
-From the crowdsourcing results of all 30 responses to each preference aspect, we observe variations in user preferences for UI widgets regarding predictability, efficiency, and explorability across the eight tasks.
-For continuous value adjustment tasks, including adjusting image lightness, saturation, and hue, many users preferred preset buttons with preview overlays for their predictability and efficiency, with reasons like “The preset button’s background allows for users to figure out the lightness relatively quickly, and when clicking on the buttons, the button updates the lightness to how it looks on the buttons, allowing for users to need minimal deduction to understand what this button does”.
-For explorability, sliders were favored for adjusting lightness and saturation.
-Participants commented that “The slider allows users to slowly go through each increment and explore what effect they have on the image.
-They can go anywhere within the range to understand the effect of saturation on the image and precisely pinpoint which exact value of saturation they want with a slider”.
-A color wheel was preferred to explore hue adjustment, with comments like “The color wheel, due to it allowing you to select exactly what you want and visualizing it, allowing you to explore and experiment very easily”.
-For discrete value selection tasks, including choosing predefined values for image lightness, saturation, hue, and watermark positions, preset buttons were favored over radio buttons and dropdown menus in all three preference aspects.
-This was primarily due to the visual cues that preset buttons provided, such as user-provided reason saying that “The preset buttons include a preview of how the saturation would look like.
-For example, -1.0 is fully gray, while 1.0 is a bright rainbow.
-This allows users who may not know what saturation means to understand what each button does; it changes how colorful the image is”.
-For tasks emphasizing visual feedback, preferences varied based on whether the task involved color or position adjustments.
-In color adjustment tasks, including changing image hue, applying fall colors, and adjusting color balance, many users preferred color wheels and color pickers for predictability and explorability, while preset buttons were preferred for efficiency, similar to continuous and discrete value adjustment tasks.
-However, for tasks requiring precise color adjustment, such as matching an object’s color to a given color, there was no clear consensus regarding predictability.
+In total, we obtained 30 responses $\times$ 3 aspects $\times$ 8 tasks = 720 user preferences.
+Figures 5 and 6 show the distribution of collected user preferences.
+Figure 5. User preference distribution of tasks 1-4 (task set 1) collected in the crowdsourcing study.
+[…]
+This figure consists of four rows, showing the crowdsourcing results of tasks 1 to 4 in each row.
+Each row has three columns, depicting the user preference distribution of the three preference aspects in each column as bar charts.
+Figure 6. User preference distribution of tasks 5-8 (task set 2) collected in the crowdsourcing study.
+[…]
+This figure consists of four rows, showing the crowdsourcing results of tasks 5 to 8 in each row.
+Each row has three columns, depicting the user preference distribution of the three preference aspects in each column as bar charts.
+5.4.2. Analysis of User Preferences
+We observe variations in user preferences for UI widgets regarding predictability, efficiency, and explorability across the eight tasks.
+Continuous value adjustment: Users preferred preset buttons with preview overlays for their predictability and efficiency.
+For explorability, sliders were favored for adjusting lightness and saturation, while a color wheel was preferred to explore hue adjustment.
+Discrete value selection: Preset buttons were generally favored over radio buttons and dropdown menus in all three preference aspects, primarily due to the visual cues provided.
+Tasks emphasizing visual feedback: User preferences varied based on whether the task involved color or position adjustments.
+(1) In color adjustment tasks emphasizing exploration, many users preferred color wheels and color pickers for predictability and explorability, while preset buttons were selected for efficiency, similar to continuous and discrete value adjustment tasks.
+(2) However, for tasks requiring precise color adjustment, such as matching an object’s color to a given color, there was no clear consensus regarding predictability.
 Participants were evenly split among text fields, color wheels, color pickers, and preset buttons.
-The reasons for their preferences included “Since I know what my target is, just typing it gives me what I want” for the text field, “It is easy to visualize the color” for the color wheel, “The color picker would accurately allow me to pinpoint the target color green and not have a doubt that it is correct” for the color picker, and “The preset button convey closely what color the rocket will have” for the preset buttons.
-Regarding efficiency, most users preferred text fields, as they could quickly “enter the hex number and see the desired result”.
+Regarding efficiency, most users preferred text fields.
 For explorability, color wheels and color pickers were the most favored.
-In position adjustment tasks, including placing a watermark or highlighting a specific area of an image, preset buttons and directly clicking on the image were preferred for predictability and efficiency, with reasons like “The presets give an indication of what to expect and wouldn’t confuse me in this task” and “Clicking is both intuitive and efficient and I know what I’$l$ get”.
-For explorability, users favored both clicking on the image since it was “Easy to just click in slightly different places as needed to get the area where I want it”, and using sliders because “You can totally visually discern where you want the watermarks and/or image to go to as it is continuous”.
-These collected user preferences are incorporated to guide UI widget reasoning in our user evaluation of the LLM-driven UI widget generation system.
-We present the user evaluation details and how the crowdsourced user preferences affect reasoned widgets for user evaluation tasks in the following section.
+(3) On the other hand, in position adjustment tasks, preset buttons and directly clicking on the image were preferred for predictability and efficiency, while for explorability, users favored both clicking on the image and using sliders.
 6. User Evaluation
-In this section, we introduce our user evaluation to assess our implemented system for the crowdsourced library-enhanced UI widget generation framework.
+In this section, we introduce the user evaluation to assess our implemented system for the CrowdGenUI framework.
 We outline the research questions to evaluate in the study, describe the evaluation setup, and present the evaluation results.
 6.1. Research Questions
-We aim to study the following research questions in our user evaluation, focusing on the effectiveness and generalizability of the UI widget generation framework and the impact of integrating the crowdsourced UI widget preference library:
-  • RQ1—Adaptivity of the framework to new, related tasks: Does the UI widget generation framework enable the generalization of crowdsourced library-enhanced widget reasoning and generation to perform tasks that differ from, but relate to, the original crowdsourcing tasks?
-  • RQ2—Impact of the library on aligning generated widgets with user preferences: Can the crowdsourced UI widget preference library, which is integrated into the UI widget generation framework, help generated widgets align more closely with user preferences for predictability, efficiency, and explorability compared to those generated without a library?
+We aim to study the following research questions in our user evaluation, focusing on the effectiveness and generalizability of the proposed framework and the impact of integrating the user preference library on UI generation:
+RQ1—Adaptivity of the framework to new, related tasks: Does the framework enable the generalization of crowdsourced library-enhanced widget reasoning and generation to perform tasks that differ from, but relate to, the original crowdsourcing tasks?
+RQ2—Impact of the library on aligning generated widgets with user preferences: Can the crowdsourced UI widget preference library help LLM-generated widgets align more closely with user preferences for predictability, efficiency, and explorability compared to those generated without a library?
 RQ3—Effect of library size on enhancing LLM reasoning: Does the size of the crowdsourced UI widget preference library affect LLM widget reasoning effectiveness?
-Which size of the library provides the strongest support for LLM reasoning capabilities?
-In the following subsections, we introduce the user evaluation setup designed to answer these research questions and analyze our evaluation results in alignment with them.
-7.2. Platform and Participants
-[…]
+Which size of the library provides the strongest support for reasoning capabilities?
+In the following subsections, we explain the rationale of these RQs, introduce the user evaluation setup designed to answer them, and analyze our evaluation results in alignment with them.
+6.2. Participants
+We conducted the user evaluation on Prolific (Prolific, 2024) and employed the same recruitment criteria as those in the crowdsourcing study to ensure consistency in participant background.
+The participants worked on the user evaluation tasks on desktops using a mouse and keyboard.
+In total, we recruited 78 participants (Age: 29.81 $\pm$ 9.09, Male: 56, Female: 19, Non-binary: 3) and compensated each participant 10 USD for the 30-minute study.
 6.3. Evaluation Setup
@@
 Full details of the user evaluation briefing and instructions can be found in Appendix C.
-Following that, they worked on three tasks and answered six questions for each task to provide their preferences for LLM-generated widgets.
+Following that, they worked on three tasks and answered six questions for each task to provide their feedback on LLM-generated widgets.
 6.3.2. User Interface
@@
 After interacting with all the offered widgets, participants were shown a preference selection panel where they chose their favored options for each task.
-(a) Task description and LLM-generated UI widgets to perform the task.
-(b) User preference selection panel for pairwise comparison. The panel contains six questions for each task, and we show one of them due to space limits.
-Figure 7. User evaluation user interface for task: adjusting image tint (image_adjust_tint).
-7.3.3. Tasks and Preference Aspects
-The user study tasks differ from the crowdsourcing tasks but still fall within the same three categories, promoting either exploration or precision (recall Section 6.2.2).
-We provide an overview of these tasks and their categories below, with detailed descriptions available in Table 2.
-These tasks are designed to address RQ1.
-Note that two tasks belong to the design domain, i.e., design_align_text and design_position_logo, expanding beyond the image editing domain emphasized in the crowdsourcing study.
-Through these user evaluation tasks, we aim to test our framework’s generalizability to new tasks in both image editing and additional domains.
+Figure 7.
+User evaluation user interface for task: adjusting image tint (image_adjust_tint).
+The preference selection panel contains six questions for each task, and we show one of them due to the space limit.
+[…]
+This figure shows an example of the user evaluation UI consisting of three components.
+From top to bottom, the components are the task description, the UI widget interaction panel to perform the task, and the user preference selection panel.
+The task description specifies the goal that the user is expected to achieve.
+Below it, the UI widget interaction panel shows multiple widgets on the left and an image to edit on the right.
+At the bottom, the preference selection panel allows users to select their preferred widget option and write their reasons for a certain preference aspect.
+6.3.3. Tasks
+The user study tasks differ from the crowdsourcing tasks but still fall within the same three categories (recall Section 5.3.2). The detailed task descriptions are in Table 2.
 Continuous value adjustment: Tasks include adjusting image exposure, tint, and temperature.
-  • Discrete value selection: Tasks include choosing predefined values for image exposure, tint, and temperature adjustment, text alignment, and logo positioning.
+Discrete value selection: Tasks include choosing predefined values for image exposure, tint, temperature adjustment, text alignment, and logo positioning.
 Tasks emphasizing visual feedback: Tasks include changing an image’s color to spring colors, aligning text with the margin, and placing a logo.
 Table 2. User evaluation task names and descriptions.
-Task Name Task Description
-1 image_adjust_exposure Adjust the image exposure by both decreasing and increasing it. Find the exposure level that appears the most natural and visually pleasing to you.
+[…]
+This table consists of two columns: task name and task description.
+In each row, the number, name, and description of each user evaluation task are detailed.
+Task Name Task Description 1 image_adjust_exposure Adjust the image exposure by both decreasing and increasing it.
+Find the exposure level that appears the most natural and visually pleasing to you.
 2 image_adjust_tint Adjust the image tint to shift its color balance, creating both subtle and dramatic effects by adding a yellow or magenta hue.
-3 image_adjust_temperature Adjust the image temperature to warm and cool tones. Experiment with different settings to achieve the most desired effect for you.
+3 image_adjust_temperature Adjust the image temperature to warm and cool tones.
+Experiment with different settings to achieve the most desired effect for you.
 4 image_change_to_spring Transform the image to reflect vibrant spring colors, adding fresh, lively hues.
-5 design_align_text Align the text “Poster Title” perfectly along one of the margins. Experiment with various positions to achieve a balanced and visually pleasing design.
+5 design_align_text Align the text “Poster Title” perfectly along one of the margins.
+Experiment with various positions to achieve a balanced and visually pleasing design.
 6 design_position_logo Experiment with various placements for the logo to determine the most visually appealing position that enhances visibility and complements the overall design of the image.
-The preference aspects: predictability, efficiency, and explorability, are drawn from the crowdsourcing study.
-We use the same aspects to evaluate whether user preferences collected from crowdsourcing can effectively lead to preference-aligned widget generation for tasks in the user evaluation.
+These tasks are designed to probe the framework’s generalizability beyond the original crowdsourcing scope to address RQ1.
+First, the tasks aim to evaluate the framework’s ability to reason UI widgets for nonoverlapping tasks involving data manipulation requirements varying from continuous and discrete value adjustments to tasks emphasizing visual feedback.
+Second, the inclusion of design tasks (e.g., text alignment and logo positioning) seeks to test cross-domain applicability, determining whether the system can extend beyond image editing despite the crowdsourced data’s focus on this domain.
+Third, the tasks are structured to balance precision with open-ended exploration, ranging from objective adjustments (e.g., margin alignment) to subjective, creativity-driven goals (e.g., change to spring colors).
+By replicating the original task categories while introducing novel domains and nuanced objectives, the study aims to evaluate if the framework can generalize to unseen tasks without domain-specific overfitting.
+6.3.4. User Preference Aspects
+The preference aspects, predictability, efficiency, and explorability, are drawn from the crowdsourcing study.
+We use the same aspects to evaluate whether user preferences collected from crowdsourcing can effectively lead to user preference-aligned widget generation for new tasks in the user evaluation.
 This setup addresses RQ2.
-7.3.4. UI Widgets for the Evaluation Tasks
-Figure 8 shows an example of LLM-reasoned widgets for task image_adjust_tint enhanced by three different sizes of the crowdsourced preference library: withlib10, withlib25, and withlib30, as well as without any library: withoutlib.
+6.3.5. UI Widgets Generated with Different Library Sizes
+Before we present the study design to address RQ3, we first explain our motivation for studying this question based on an analysis of LLM-generated UI widgets.
+Figure 8 shows an example of LLM-reasoned widgets for task image_adjust_tint enhanced by three different sizes of the crowdsourced preference library as well as without any library—withoutlib.
 Withlib10, withlib25, and withlib30 stand for the crowdsourced library with 10, 25, and 30 unique responses to each preference aspect.
-As introduced in Section 6.3.1, we obtain 30 responses to each aspect in total, i.e., withlib30, and use its subsets to construct withlib10 and withlib25.
-In Figure 8, the percentages in the pie charts show the frequency of each type of UI widget reasoned by the LLM over 10 iterations.
-For instance, in the pie chart for efficiency reasoned by withoutlib (the rightmost chart in the second row), the preset buttons are reasoned 8 times (80%), while the slider is reasoned 2 times (20%).
-Reasoned widgets differing in multiple iterations are caused by the ad-hoc nature of LLMs, i.e., they do not always generate all possible widgets in a single pass, similar to the observations in previous work (Barke et al., 2023; Vaithilingam et al., 2024).
-This observation reveals varying consistencies of LLM-reasoned widgets with different library sizes.
-Thus, we leverage the reasoning frequency to indicate the recommendation strength for each type of widget by assigning each a score in our user evaluation.
-For example, the frequencies of 8 for the preset buttons and 2 for the slider are used as scores 8 and 2 for them, as seen in Figure 7(b) Option 2.
-Due to space limit, we provide the LLM-reasoned widgets with different library sizes for all six tasks in Appendx A.
-As seen in Figure 8, LLM-reasoned UI widgets differ significantly with and without the crowdsourced libraries across the three preference aspects.
-For predictability, the color wheel is the most frequently reasoned widget when using the library, while the slider takes precedence in the absence of a library.
-In terms of efficiency, both withlib10 and withlib30 favor the color picker, while withlib25 most frequently reasons the color wheel, followed closely by the preset buttons and color picker.
-Without a library, preset buttons are the most commonly reasoned widgets.
-Regarding explorability, the color wheel is reasoned the most when libraries are utilized, whereas the color picker is the top choice without a library.
-Similar reasoning differences can be found in the reasoned widgets for other tasks in Appendx A.
-These findings prompt us to address RQ3: What is the optimal size of the crowdsourced library that maximizes LLM reasoning capabilities preferred by users with regard to the three preference aspects?
-To investigate this question, we conduct pairwise comparisons between libraries and collect user reasons for their choice.
+As introduced in Section 5.4, we obtain 30 responses to each aspect in total and use its subsets to construct withlib10 and withlib25.
+Due to space limit, we provide the LLM-reasoned widgets with different library sizes for all six tasks in Appendix A.
 Figure 8.
@@
 Withlib10, withlib25, and withlib30 refer to using the crowdsourced libraries with 10, 25, and 30 user responses for all aspects and tasks, while withoutlib refers to using no library.
+[…]
+This figure consists of three rows, showing the LLM-reasoned widgets for the three preference aspects in each row.
+Each row has four columns, depicting the reasoned widget frequency as percentages using the four library sizes—withlib10, withlib25, withlib30, withoutlib—in each column as bar charts.
+As seen in Figure 8 (and figures in Appendix A), the percentages in the bar charts show the frequency of each type of UI widget reasoned by the LLM over 10 iterations.
+For instance, in the bar chart for efficiency reasoned by withoutlib (the rightmost chart in the second row of Figure 8), the preset buttons are reasoned 8 times (80%), while the slider is reasoned 2 times (20%).
+Reasoned widgets differing in multiple iterations are caused by the ad-hoc nature of LLMs (Barke et al., 2023; Vaithilingam et al., 2024), i.e., they do not always generate all possible widgets in a single pass.
+This variability is not merely a limitation—it reflects the LLM’s capacity to explore diverse design possibilities through probabilistic reasoning (Ozturkler et al., 2022), producing a dynamic space of UI designs.
+By contrast, fixed UIs with built-in widget suggestions rely on static, preprogrammed heuristics that prioritize generality.
+While such systems may use common user preferences for UI suggestion, they limit broader possibilities by enforcing a one-size-fits-all UI design approach, which struggles to accommodate niche workflows or evolving user expectations (Findlater and Gajos, 2009).
+Instead, LLM’s dynamic, iterative UI generation surfaces a spectrum of context-aware widgets, transforming variability into a mechanism for adaptive UI suggestions.
+Thus, to leverage the design variability enabled by the non-deterministic LLM outputs, we use the reasoning frequency to indicate the recommendation strength for each type of widget by assigning each a score in our user evaluation.
+For example, the frequencies of 8 for the preset buttons and 2 for the slider are used as scores 8 and 2 for them, as seen in Figure 7 Option 2.
+Aggregating widget frequencies across iterations acts as an implicit voting system, where higher-frequency widgets signal stronger consensus of user preference, akin to collaborative filtering in recommender systems (Schafer et al., 2007).
+This emergent prioritization reduces the manual effort required to curate or update UI suggestions, as the LLM dynamically reweights suggestions based on reasoned relevance.
+The scoring of widgets by frequency allows users to quickly identify both common and edge-case tools, mirroring how human designers iteratively refine interfaces through trial and error of multiple possibilities (Nielsen, 2002) and how recommender systems balance relevance with serendipity when presenting content (De Gemmis et al., 2015; Kotkov et al., 2016).
+As we observe that the types and frequencies of LLM-reasoned UI widgets vary depending on the size of the library, we are motivated to explore RQ3—the optimal size of the library that maximizes LLM’s UI reasoning capabilities.
+This question is crucial not only for our study but also for broader machine learning research, where determining the optimal dataset size is essential for balancing cost and performance.
+A well-sized library should be large enough to represent diverse user inputs, but cannot grow indefinitely due to redundancy and cost.
+Moreover, overly large libraries risk overfitting the model or diluting meaningful patterns in user preference, making it crucial to find a size that is both effective and efficient.
+To this end, we perform pairwise comparisons of widgets generated with libraries of varying sizes, which are introduced in the following subsection.
 6.3.6. Pairwise Comparisons
-We conduct pairwise comparisons for the reasoned widgets, their reasoned frequency, and the LLM-provided reasoning enhanced by withlib10, withlib25, withlib30, and withoutlib.
-By including widgets reasoned over multiple iterations, 10 in our study, we aim to include as many relevant widgets as possible in the evaluation, enabling a comprehensive comparison of LLM reasoning capabilities augmented by different sizes of the libraries.
-Participants need to select their preferences by comprehensively considering the widgets, their frequencies, where higher frequency indicates a stronger recommendation, and the reasoning, which explains why each widget is deemed appropriate for the task and preference aspect (recall Figure 7(b) for an example of pairwise preference selection).
-Each preference aspect in each task results in six pairwise comparisons, as outlined in Table 3.
+We conduct pairwise comparisons for LLM-generated UI widgets when withlib10, withlib25, withlib30, and withoutlib are used.
+Participants need to select their preferences by comprehensively considering the widgets and their frequencies, where a higher frequency indicates a stronger recommendation, and the LLM’s reasoning, which explains why each widget is deemed appropriate for the task and preference aspect.
+Consequently, each preference aspect in each task results in six pairwise comparisons, as outlined in Table 3.
 Table 3. Pairwise comparisons of different crowdsourced library sizes for each preference aspect in each task. Pair orders are randomized in the user evaluation.
-Library Sizes Compared
-1 withlib10 vs. withlib25
-2 withlib10 vs. withlib30
-3 withlib10 vs. withoutlib
-4 withlib25 vs. withlib30
-5 withlib25 vs. withoutlib
-6 withlib30 vs. withoutlib
-In total, this gives us 6 pairs $\times$× 3 aspects $\times$× 6 tasks = 108 comparisons.
+[…]
+This table details the six possible pairwise comparisons of library sizes.
+Library Size Comparison 1 withlib10 vs.
+withlib25 2 withlib10 vs.
+withlib30 3 withlib10 vs.
+withoutlib 4 withlib25 vs.
+withlib30 5 withlib25 vs.
+withoutlib 6 withlib30 vs.
+withoutlib
+In total, this gives us 6 pairs $\times$ 3 aspects $\times$ 6 tasks = 108 comparisons.
 To prevent user fatigue, each participant was assigned 18 pairs, which consisted of all six pairs for each aspect across three tasks.
@@
 Each participant provided preferences for all three aspects of their assigned task set.
-The tasks were arranged in all six possible order combinations, and the aspect order was balanced using the Latin Square method (Colbourn, 2010).
+The tasks were presented in all six possible permutations to account for order effects, and the sequence of aspects was counterbalanced using the Latin Square method (Colbourn, 2010).
 See Table 4 for the detailed study setup.
-The six pairwise comparisons for each aspect of each task were presented in random order.
-To ensure robust statistical results, each pairwise comparison was tested by 12 users, resulting in 6 pairs $\times$× 12 users = 72 total participants.
 Table 4.
@@
 The six user evaluation tasks are divided into two task sets.
-In each set, we arrange all six order combinations of tasks and aspects and randomize orders of pairwise comparisons.
-Task Set 1 Task Set 2
-Order Task Name Aspect Order Task Name Aspect
-6 $\times$× image_adjust_exposure Predictability 6 $\times$× image_adjust_tint Predictability
-image_adjust_temperature Efficiency image_change_to_spring Efficiency
-design_align_text Explorability design_position_logo Explorability
-6 $\times$× image_adjust_exposure Efficiency 6 $\times$× image_adjust_tint Efficiency
-image_adjust_temperature Explorability image_change_to_spring Explorability
-design_align_text Predictability design_position_logo Predictability
-6 $\times$× image_adjust_exposure Explorability 6 $\times$× image_adjust_tint Explorability
-image_adjust_temperature Predictability image_change_to_spring Predictability
-design_align_text Efficiency design_position_logo Efficiency
+Within each set, all six possible permutations of task and aspect orders are employed to control for order effects, and the sequence of pairwise comparisons is randomized.
+[…]
+This table consists of two components: the evaluation setup for task sets 1 and 2.
+In each component, the sub-table contains three columns: the total order number of tasks, task names, and preference aspects.
+Task Set 1 Task Set 2 Order Task Name Aspect Order Task Name Aspect 6 $\times$ image_adjust_exposure Predictability 6 $\times$ image_adjust_tint Predictability image_adjust_temperature Efficiency image_change_to_spring Efficiency design_align_text Explorability design_position_logo Explorability 6 $\times$ image_adjust_exposure Efficiency 6 $\times$ image_adjust_tint Efficiency image_adjust_temperature Explorability image_change_to_spring Explorability design_align_text Predictability design_position_logo Predictability 6 $\times$ image_adjust_exposure Explorability 6 $\times$ image_adjust_tint Explorability image_adjust_temperature Predictability image_change_to_spring Predictability design_align_text Efficiency design_position_logo Efficiency
+6.3.7. Data Analysis
+We employ the Chi-squared ($\chi^{2}$) test (Pearson, 1900) to analyze the pairwise comparisons.
+The $\chi^{2}$ test evaluates whether observed preference selection frequencies deviate significantly from expected frequencies under the null hypothesis of no association between library size and user preference.
+We implement this analysis programmatically using SciPy (SciPy, 2024) with significance levels denoted by asterisks: * for $p<0.05$, **∗ ∗ for $p<0.01$, and ***∗ ∗ ∗ for $p<0.001$.
 6.4. Results
-7.4.1. Results of Each Aspect across All Tasks
-We present the user evaluation results of each preference aspect across all tasks in Figure 9.
-As can be seen, for predictability and explorability, users preferred the LLM-reasoned widgets with libraries over those without any library.
-In terms of efficiency, LLM-reasoned widgets with libraries were slightly more favored than those without.
-When the libraries were incorporated, withlib30 was more preferred than withlib10 and withlib25 for predictability and explorability, while for efficiency, withlib25 and withlib30 were almost equally favored and more preferred than withlib10.
-Figure 9. Total user preference selection counts of three preference aspects across all tasks.
-7.4.2. Results of Each Aspect in Each Task
-Next, we examine user preferences for LLM-reasoned widgets per aspect and task.
-Figures 10 and 11 depict results of the task sets 1 and 2.
-As observed, user preferences show varying trends depending on the task.
+6.4.1. Overview of All Tasks
+We present the user evaluation results of each preference aspect across all tasks in Figures 9(a) and 9(b).
+For all three aspects, users preferred the LLM-reasoned widgets with libraries over those without any library.
+When the libraries were incorporated, withlib30 was consistently more preferred than withlib10 and withlib25.
+(a) Total selection counts.
+(b) Distribution of total selection counts. The box plot elements show the median, quartiles (25th and 75th percentiles), and the data range.
+Figure 9. Number and distribution of user preference selection counts of three preference aspects across all tasks.
+[…]
+This figure contains two sub-figures: (a) shows the total selection counts as bar charts; (b) shows the distribution of total selection counts as violin plots.
+Both sub-figures consist of three columns that present the results of the three preference aspects.
+6.4.2. Inspect Each Task
+We further present the results of each evaluated task. According to the results, user preferences show varying trends depending on the task.
 Figure 10. User preference selection counts of tasks in task set 1.
+[…]
+This figure contains three sub-figures in three rows, showing the total selection counts as bar charts for tasks in task set 1.
+All sub-figures consist of three columns that present the results of the three preference aspects.
 Figure 11. User preference selection counts of tasks in task set 2.
-In task set 1, for image_adjust_exposure (see Figure 13 for LLM-reasoned widgets), widgets reasoned with libraries were more preferred than those without a library for both predictability and explorability, with withlib25 and withlib30 being nearly equally favored.
+[…]
+This figure contains three sub-figures in three rows, showing the total selection counts as bar charts for tasks in task set 2.
+All sub-figures consist of three columns that present the results of the three preference aspects.
+Figure 10—task set 1:
+  • image_adjust_exposure (see Figure 12 for LLM-reasoned widgets): Widgets reasoned with libraries were more preferred than those without a library for both predictability and explorability, with withlib25 and withlib30 being nearly equally favored.
 For efficiency, withlib25 and withlib30 were preferred over withoutlib, and withoutlib was favored over withlib10.
-When LLM-reasoned widget types were similar by different libraries, such as 60% preset buttons and 40% slider by withlib25, and 70% preset buttons and 30% slider by withlib30, participants’ feedback suggested that the slight score difference and LLM-provided reasoning helped guide their preference selection.
-For instance, participants chose withlib30 due to “good reasonings for any user which would prefer an easier time with Preset Buttons of higher score”.
-For image_adjust_temperature (see Figure 15 for LLM-reasoned widgets), withoutlib was preferred for predictability and almost equally favored as withlib25 and withlib30 for efficiency.
-Participants who preferred sliders reasoned by withoutlib over the color wheel and color picker reasoned with the libraries highlighted the slider’s predictability.
-For example, they commented, “the slider is the most instantly understandable widget… the color wheel and color picker require a bit of deduction from the user on how to utilize them properly”.
-They also remarked on sliders’ sensitivity to value changes, such as “I like the sensibility of sliders in general, and I feel they are a very efficient means of getting a good result by eye, just by dragging your mouse left or right, so this is the best”.
-However, widgets reasoned with libraries were favored for explorability.
-For instance, participants preferred the color picker reasoned by withlib10 because “you are given 3 RGB values each ranging from 0 to 255, giving you 255^3 possibilities”, as opposed to the slider reasoned by withoutlib since “you are only given 100 options as the slider moves in intervals of 0.01”.
-With regard to design_align_text (see Figure 17 for LLM-reasoned widgets), widgets reasoned with libraries were again more favored for predictability, and trends similar to image_adjust_temperature were observed for efficiency and explorability.
-In task set 2, for image_adjust_tint (see Figure 8 or 14 for LLM-reasoned widgets), widgets reasoned with libraries remained more favored for predictability compared to those without a library.
-For efficiency, withoutlib was the most preferred option, indicating that the preset buttons were favored over the color picker and color wheel reasoned with libraries.
-Regarding explorability, withlib10, withlib25, and withoutlib were nearly equally favored, showing that participants preferred the color wheel and color picker for exploration over sliders reasoned by withlib30.
-In terms of image_change_to_spring (see Figure 16 for LLM-reasoned widgets), withlib25 and withlib30 were almost equally preferred, both more favored than withlib10 and withoutlib for predictability.
-In terms of efficiency, withlib10 and withlib25 were preferred over withlib30 and withoutlib.
-For explorability, widgets reasoned with libraries were consistently more favored than those without any library.
-Regarding design_position_logo (see Figure 18 for LLM-reasoned widgets), widgets reasoned with libraries were more favored across all three aspects.
-Withlib10 was the most preferred for predictability and efficiency, while withlib30 was the most favored for explorability.
-For predictability and efficiency, withlib10 reasoned similar widgets (click on image) to those reasoned by other libraries, but it was preferred over withlib25 and withlib30 largely due to its reasoning.
-For instance, participants commented, “Option 1 [withlib10] made more sense because it said how it takes minimal effort, instant feedback, and it’s quick and precise.
-Whereas option 2 [withlib30] said it’s just simple and takes minimal effort”.
-7.4.3. Reflection on Research Questions
-For RQ1, the results reveal that the user preferences collected from crowdsourcing can effectively generalize to tasks beyond the original crowdsourcing tasks by enhancing the LLM’s reasoning capabilities.
-Broadly speaking, across all user evaluation tasks, widgets reasoned with the library are largely favored over those reasoned without a library.
+  • image_adjust_temperature (see Figure 14 for LLM-reasoned widgets): Withlib30 was the most favored and almost equally favored as withoutlib for predictability and withlib25 for efficiency.
+In addition, widgets reasoned with libraries were consistently favored for explorability.
+  • design_align_text (see Figure 16 for LLM-reasoned widgets): Widgets reasoned with libraries were consistently more favored for all three aspects.
+Figure 11—task set 2:
+  • image_adjust_tint (see Figure 8 for LLM-reasoned widgets): Widgets reasoned with libraries were more favored for predictability compared to those without a library.
+For efficiency, withlib30 and withoutlib were almost equally preferred.
+Regarding explorability, withlib25 was the most favored.
+  • image_change_to_spring (see Figure 15 for LLM-reasoned widgets): Withlib30 was the most preferred for predictability and explorability.
+In terms of efficiency, withlib25 was preferred over the other libraries.
+  • design_position_logo (see Figure 17 for LLM-reasoned widgets): Widgets reasoned with libraries were significantly more favored across all three aspects.
+6.5. Findings
+6.5.1. Generalization to Nonoverlapping Tasks and Broader Domains [RQ1]
+The results reveal that the user preferences collected from crowdsourcing can effectively generalize to tasks beyond the original crowdsourcing tasks by enhancing LLM reasoning.
+Across all user evaluation tasks, including the tasks not in the image editing domain, widgets reasoned with the library are largely favored over those reasoned without a library.
 This highlights the adaptability and relevance of user preferences captured through crowdsourcing, suggesting that these insights can serve as a foundation to inform widget reasoning across a range of related tasks.
-While the overall trend favors library-informed reasoning, some exceptions emerge for specific tasks.
-For instance, for predictability in image_adjust_temperature and both efficiency and explorability in image_adjust_tint, widgets reasoned without the library are preferred.
-These results suggest that certain tasks may require additional consideration in the reasoning process.
-For RQ2, the results demonstrate that incorporating the crowdsourced library enhances the ability of the LLM to reason widgets that align with user preferences regarding predictability, efficiency, and explorability.
-In terms of predictability, users’ preferences captured from the crowdsourcing study lean towards preset buttons for predefined value selection, color wheels, and color pickers for color adjustments, and click on image for position changes (recall Section 6.3.2).
-These preferences have directly influenced the widgets reasoned for the user evaluation tasks, where widgets reasoned with the library adhered more closely to these favored widget types.
+Takeaway: Crowdsourced user preferences effectively enhance LLM widget reasoning, not only being favored across image editing tasks but also generalizing to inform widget reasoning for relevant tasks beyond image editing.
+6.5.2. Preference Alignment of Generated UI Enhanced by Crowdsourced Library [RQ2]
+The results demonstrate that incorporating the crowdsourced library enhances the LLM’s ability to reason widgets that align with user preferences.
+For predictability, users’ preferences captured from the crowdsourcing lean towards preset buttons for predefined value selection, color wheels, and color pickers for color adjustments, and click on image for position changes (recall Section 5.4).
+These preferences have directly influenced the widgets reasoned for the user evaluation tasks, where widgets reasoned with the library adhered closely to these favored widget types.
 In contrast, when the library is not used, the LLM reasons sliders for most tasks regardless of the task categories.
 For efficiency, similar patterns emerge—preset buttons, color wheels, color pickers, and click on image are repeatedly preferred for their ease and speed in achieving task goals.
-The LLM, when equipped with the crowdsourced library, is able to reason toward these efficient widgets, whereas, without the library, the reasoning leans more heavily on sliders and preset buttons for most tasks without too much differentiation for task categories.
+The LLM, when equipped with the crowdsourced library, can reason toward these efficient widgets, whereas, without the library, the reasoning leans more heavily on sliders and preset buttons for most tasks without too much differentiation for task categories.
 Regarding explorability, the user preference trends show that sliders are favored for continuous value adjustments, color wheels are preferred for color changes, and click on image is liked for position adjustments.
 Without the library, however, the LLM tends to reason sliders to allow exploration for continuous value adjustments, color pickers for color changes, and sliders and text fields for position adjustments.
-The incorporation of the crowdsourced library enables the LLM to capture the need for exploration in tasks like color change and position adjustment by effectively aligning the widget choices with user preferences.
-For RQ3, we observe that withlib30, containing the most crowdsourced data, emerges as the most preferred library across participants, particularly for predictability and explorability, while performing almost equally with withlib25 in terms of efficiency.
-This pattern, supported by the distribution of user preferences across all tasks, as shown in Figure 12, highlights withlib30 as the most reliable for all three aspects due to the richness of its crowdsourced data.
-This consistency of user preferences underscores the value of extensive, diverse data in reasoning well-aligned widget options for users.
-On the other hand, withoutlib shows more significant variability in user preferences, particularly in predictability and efficiency, suggesting that when no library is used, LLM reasoning lacks the consistency that is otherwise afforded by crowdsourced data.
-While withlib10 and withlib25 are moderately favored for predictability and efficiency, they do not perform as well in explorability as withlib30, further emphasizing the benefit of larger datasets in supporting widget reasoning for tasks requiring exploration.
-However, examining each task individually reveals significant variation in preferred libraries for different preference aspects.
-This indicates that while larger libraries lead to consistent preference-aligned widget reasoning across multiple tasks in general, for individual tasks, the utility of LLM-reasoned widgets for specific preference aspects may need additional evaluation.
-Figure 12. Distribution of user preference selection counts of three preference aspects across all tasks.
+The incorporation of the crowdsourced library enables the LLM to capture the need for exploration in tasks like color and position adjustment by aligning the widget choices with user-preferred ones.
+Takeaway: Integrating the crowdsourced library enables the LLM to reason widgets that more accurately align with user preferences for predictability, efficiency, and explorability, while its absence leads to more generic and less tailored reasoning.
+6.5.3. Effect of Library Sizes on UI Generation [RQ3]
+Both the total selection counts and their distributions, shown in Figures 9(a) and 9(b), highlight withlib30 as the most reliable for all three aspects due to the richness of its crowdsourced data.
+This consistency of user preferences underscores the value of extensive, diverse data in reasoning user-aligned widgets.
+On the other hand, withoutlib shows significant variability in user preferences, suggesting that when no library is used, LLM reasoning lacks the consistency that is otherwise afforded by crowdsourced data.
+However, examining each task individually reveals some variation in preferred libraries for different preference aspects.
+This indicates that while larger libraries lead to consistent preference-aligned widget reasoning in general, tasks like image_adjust_temperature and image_adjust_tint show that withoutlib is almost equally favored as withlib30 regarding certain aspects and may need additional evaluation.
+Takeaway: Extensive crowdsourced data makes it the most reliable and consistent for widget reasoning across tasks, while libraries with fewer data lead to greater variability in user preferences.
 7. Discussion
-Based on the analysis of user evaluation results, this section discusses the implications of these results from three perspectives: the balance between task-generic and task-specific LLM widget generation, the alignment with user preferences for user-centered UI design, and the use of user preference data to create UIs that adapt to personalized use cases and evolve over time.
-8.1. Generalizability of User Preference Data to Support Various Tasks
-The user evaluation findings suggest that crowdsourced UI widget preference libraries provide an effective foundation to improve LLM reasoning across tasks beyond the original scope of crowdsourcing.
-From the tasks tested in the user evaluation, we have shown the task generalizability not only in the image editing domain but also in the design domain (recall tasks 5 and 6 in Table 2).
-This generalization ability can reduce the need for task-specific widget customization using automatic LLM widget generation, streamlining the UI design process.
-To further enhance this generalizability, the scope of crowdsourced tasks can be expanded by incorporating a more comprehensive array of task categories beyond continuous value adjustment, discrete value selection, and color and position changes.
-In addition, we can deepen the exploration within specific task categories to involve more complex aspects of data manipulation.
-For instance, our study emphasized tasks that required exploration and precision within the three task categories, and this can be extended to tasks on content editing or value adjustment that emphasize simplicity, speed, or flexibility.
-Also, while we focused on single-value adjustment, this setup can be extended to multi-dimensional data manipulation or compound parameter adjustments.
-By carefully curating the tasks for crowdsourcing, we envision that LLMs can be equipped with crowdsourced insights to reason UI widgets in both task-generic and task-specific contexts, adapting to diverse user needs and task scenarios.
-Moreover, combining task-generic and task-specific UI widget reasoning could provide a solution for cases where generalized user preferences alone are insufficient.
-In our user evaluation, preference aspects of a few tasks showed that user preferences were better supported without the library (recall Sections 7.4.2 and 7.4.3).
-These findings highlight that crowdsourced preferences do not universally apply to all tasks, and some require more refined, context-sensitive reasoning.
-To address this, an adaptive approach that integrates task-specific customization with the general crowdsourced library can potentially better accommodate unique user interaction demands to ensure the LLM delivers more tailored and effective UI widget suggestions across a wide range of tasks.
-For example, for the efficiency aspect of image_adjust_tint, users showed a preference for preset buttons generated by withoutlib over the color wheel and color picker reasoned with the library.
-Incorporating this insight into the LLM reasoning process by assigning higher weights to preset buttons for tasks involving image tint adjustments and related operations can lead to more preference-aligned UI generation than directly using the library.
-This approach leverages the flexibility of LLM-based widget generation by taking updated user insights to adapt to both generic and specific task requirements.
-8.2. Towards User-Centered UI Widget Selection and Design
-The crowdsourcing results reveal significant variation in user preferences for UI widgets based on different data manipulation needs, emphasizing the importance of a user-centered approach to UI widget selection.
-This approach contrasts with the traditional designer-dominant method, where UI widgets are often pre-selected by software designers.
-For instance, tools like Photoshop provide a slider and a text field for adjusting image lightness, saturation, and hue, which offer a one-size-fits-all solution.
-However, our crowdsourced data shows that users prefer distinct widgets for various data manipulation requirements, such as preset buttons for efficient adjustment of lightness and saturation and a color wheel to explore multiple hue options (recall Section 6.3.2).
-This divergence reveals a gap between user-preferred and designer-selected UI widgets.
-This gap highlights that static, predefined UI widgets in most software may not always meet the diverse needs of users.
-LLM-supported UI generation offers a promising solution to bridge this gap by generating dynamic widgets.
-With the crowdsourced preference library, we have extended the existing LLMs’ ability to reuse user-preferred widgets based on user preference insights and, therefore, generate preference-aligned widgets for a broad range of tasks.
-This framework can be used to enable the creation of user-centered interfaces that align with the preferences collected from users to augment the traditional UI widget selection method that heavily relies on designers.
-Furthermore, the user-centered UI design approach can be expanded to accommodate a broader spectrum of preference aspects and cater to more diverse user groups.
-Our study focuses on three aspects of user preferences, demonstrating the prototype’s effectiveness in guiding LLM reasoning, but future research could incorporate additional dimensions of UI design.
-For example, prior work on UI design principles (McKay, 2013; Lidwell et al., 2010; Chen et al., 2021), which covers a number of guidelines, can be used to inform crowdsourcing efforts.
-Among them, accessibility is a particularly important dimension that can be addressed by collecting user preferences from individuals with disabilities.
-Crowdsourced data from these users could inform the design of accessible UI widgets, ensuring that interfaces meet their specific needs.
-This method would not only aid LLMs in generating accessible widgets that benefit users but also provide valuable insights for software designers.
-For example, LLMs can be equipped with crowdsourced data to suggest widget sizes optimized for low-vision users or alternative interaction techniques for those with motor impairments.
-They can also be prompted to offer reasonings for their suggestions, which could serve as references for design rationale to assist designers.
-Thus, the LLM-reasoned widgets can potentially augment the ideas proposed in prior research, such as SUPPLE (Gajos and Weld, 2004) and SUPPLE++ (Gajos et al., 2007), to develop dynamic UIs that accommodate motor and vision capabilities.
-8.3. Opportunities for Adaptive UIs
-The comparison between different library sizes reveals that larger libraries tend to produce more consistent preference-aligned widget reasoning across various tasks; meanwhile, it highlights the potential of smaller libraries for cost-effective and adaptive UI design.
-Our study shows that even with a library of just 10 user responses, i.e., withlib10, key insights into user preferences can be effectively captured.
-In some tasks and aspects, such as the predictability of adjusting image exposure and tint, the efficiency of changing colors to spring tones and positioning a logo, and the explorability of aligning text and adjusting tint, smaller libraries performed as well or even better than larger ones.
-Small-scale libraries are quick to collect and easy to evolve, providing efficient UI design suggestions for designers to reference in the design process.
-Even at the extreme, i.e., using no library at all, LLMs can reason UI widgets that, in some instances, outperformed those based on larger libraries.
-This demonstrates that LLMs can aid in UI widget reasoning and selection without requiring extensive human input, which offers designers a low-cost way to access user preferences during UI design.
-Such adaptability of small libraries can significantly reduce the time and resources required to gather large datasets while still yielding meaningful user-centered insights.
-The adaptability enabled by small-scale libraries for UI design can be extended to deeper UI personalization, which reflects individual preferences and supports long-term UI evolution.
-By collecting a few samples from individual users, these libraries can be tailored to represent personalized preferences for UI widget generation.
-Such personalization can allow users to build workflows that align with their specific needs.
-Moreover, personalized libraries have the potential to evolve alongside users as their skills and expertise grow.
-For example, the LLM could initially generate simple, beginner-friendly widgets and, over time, adapt to offer more advanced, feature-rich widgets as the user becomes more proficient by using the user’s updated personal library it uses for widget reasoning.
-This dynamic evolution enables users to progressively engage with more complex tools to support an adaptive and long-term learning curve.
-This kind of UI evolution not only streamlines immediate tasks but also fosters continuous improvement in tool use, helping users to gradually master more sophisticated functions without overwhelming them at the start.
+In this section, we reflect on our findings and discuss their implications for crowdsourcing-enhanced human-centered UI generation.
+7.1. Improving Task Generalizability with Curated Crowdsourcing Setup
+Importance of task curation to collect helpful user data that can boost LLM-based UI generation: The user evaluation findings indicate that the crowdsourced user preference library improved the LLM’s reasoning capabilities, extending the library’s effectiveness beyond its initial intended task scope and domain.
+Such generalizability highlights the user data’s potential to serve as a foundational element to guide the creation of versatile, human-centered UIs across diverse domains through careful task curation.
+This can contribute to a significant reduction in the necessity for task-specific UI customizations.
+For instance, if we collect user-preferred UIs to adjust image tones, this insight can be helpful in indicating user preferences for a range of similar tasks involving continuous color adjustment.
+In our implementation, the effectiveness of incorporating tasks belonging to the three categories involving diverse data manipulation requirements (recall Section 5.3.2) has been demonstrated through our evaluation.
+This implies that through well-curated task design, UI designers can collaborate with LLMs that are equipped with user data to prototype designs to meet users’ expectations with reduced additional user testing.
+Extending the breadth and depth of task categories: To further enhance the user data’s generalizability, future efforts can strategically encompass a broader spectrum of task categories.
+In addition to this work’s focus on image parameter adjustment, other potential task categories include audio, video, data visualization editing, and others introduced in Section 2.
+These domains often involve distinct interaction styles.
+For instance, video editing might require precise temporal controls like scrubbers and frame-level sliders, while data visualizations might prefer touch-based navigation on tablets versus hover-driven interactions on desktops.
+Because each domain brings unique data manipulation, their UI designs may not be easily generalizable across domains.
+As such, multi-domain user data collection is essential to ensure UI recommendations remain relevant and effective for diverse task contexts.
+On the other hand, future research can also aim for a deeper exploration of specific task categories, focusing on more complex and intricate forms of data manipulation.
+For instance, for image editing, we can extend the scope from single-value to multi-dimensional or compound parameter manipulations.
+This is a common need as image editing often consists of a combination of multiple parameters, e.g., if a user wants to correct a low-exposure image, they may increase both image saturation and brightness.
+Thus, a dataset containing user data on complex or open-ended tasks may help inform the LLM-based UI generation framework to create a widget that combines multi-parameter adjustment or to generate widget combinations.
+Investigating extended task complexities can help uncover the dependence of specific parameters and the relevance of high-level task requirements, thus refining the granularity of crowdsourced data to bolster the applicability of LLM-generated UIs in addressing complicated task contexts.
+7.2. Designing User-Focused UIs with Insights Collected from Target Audience
+Tailoring UIs to target users favored over traditional fixed UI designs: Our crowdsourcing results highlight variations in user preferences for UI widgets, underscoring the need for a user-centered approach to widget selection.
+This contrasts with traditional ways of UI design by most software and applications, i.e., UIs are fixed and consist of designers’ pre-selected UI elements.
+For instance, Photoshop offers a slider and text field for adjusting image lightness, saturation, and hue.
+However, our crowdsourced data reveals various user-preferred widgets that go beyond these given options.
+This divergence exposes a gap that static, designer-selected widgets sometimes fail to meet diverse user needs.
+To this end, our framework utilizes LLM-based dynamic UI design, which generates UIs aligned with user preferences and task contexts.
+This framework facilitates the creation of user-centered interfaces with direct end-user input versus the traditional designer-driven methods.
+Collecting specific, targeted user data to guide tailored UI design: Our user-centered framework can further evolve to encompass a broader range of preference dimensions and serve more diverse user populations.
+While this work uses three preference aspects to demonstrate the framework’s efficacy, future work can explore additional dimensions by using established UI design principles (McKay, 2013; Lidwell et al., 2010; Chen et al., 2021; Shneiderman, 2000; Norman, 1983).
+Among them, we take accessibility as an example to show how other dimensions can be incorporated.
+Accessible UI design is a crucial dimension and is addressable through crowdsourcing preferences from users with disabilities.
+The collected data would enable LLMs to reason accessibility considerations directly based on the target user’s data.
+For instance, users with visual impairments might prefer high-contrast color schemes, screen reader-friendly labels, or voice-based navigation.
+Those with motor impairments may favor interfaces with larger clickable targets, simplified gesture controls, or customizable input methods based on varied motor abilities (Gajos and Weld, 2004; Gajos et al., 2007).
+These accessibility preferences, once incorporated into the user library, can be used to guide LLM reasoning during UI generation.
+In turn, the LLM can recommend design adaptations tailored to specific user groups, such as suggesting larger widget sizes or reducing the need for precise cursor movements, and provide explanations grounded in the collected data.
+This not only supports more inclusive UI outcomes for end users but also makes the rationale behind design recommendations transparent for designers’ reference.
+7.3. Building Adaptive UIs with Scalable User Data
+Multi-scale user data to guide LLM reasoning on UI designs: Based on the comparisons of libraries of varying sizes, our results reveal that larger libraries generally yield consistent preference-aligned widget reasoning across diverse tasks.
+On the other hand, smaller libraries stand out for their cost-effectiveness and adaptability to guide UI design.
+Our study demonstrates that a library with as few as 10 user responses can effectively capture insights into user preferences.
+In specific tasks, such as adjusting image exposure, shifting colors to spring tones, and positioning logos, small-scale libraries perform comparably with their larger counterparts for certain preference aspects.
+These libraries are quick to compile, offering users and designers fast UI design recommendations without the need for extensive data collection and user testing.
+This highlights how LLMs can support UI design with cost-efficient human input for user-centered design.
+Low-cost, small-scale user data for UI personalization and evolution: The adaptability of small-scale libraries can enable deeper personalization and long-term UI evolution.
+By incorporating a few user-specific samples, the small data can be used to reflect individual preferences and tailor workflows that align with unique needs.
+Such UI personalization and evolution can be especially beneficial for beginner users who want to use professional, feature-rich software without feeling overwhelmed at the start.
+As users’ skills and expertise develop, their personal data can evolve and dynamically adjust the UIs to be more advanced over time.
+For instance, we can use our framework to initially generate simple, beginner-friendly UIs, then progressively adapt to offer sophisticated UIs with advanced tools as the user’s proficiency grows, guided by an updated personal library covering user traces.
+This evolution monitors user growth and fosters a gradual and personalized learning curve, allowing users to engage with increasingly complex tools at their own pace.
+With a similar approach, creators can also use UI personalization and evolution to tailor UIs to individual workflows.
+Although existing software allows users to show/hide tools for UI personalization (Adobe, 2024c; Blender, 2024), this often does not allow transformative UI redesign like regrouping tools into a user-created panel.
+Instead, using AI to generate UIs can programmatically create tools with personalized, evolving functionality that is tailored to creators’ own needs.
+This dynamic approach can be beneficial to enhance productivity as users are not required to learn and adapt to given UIs, but instead, UIs are tailored to their workflows.
+7.4. Exploring UI Design in Human-Centered Agentic AI
+Indispensable role of UIs in agentic AI: As agentic AI increasingly automates complex tasks, e.g., building agents with Magentic-One (Fourney et al., 2024) to write and execute code, we envision agentic AI systems to help delegate tasks that can significantly foster human productivity.
+In the meantime, alongside the pursuit of fully autonomous agentic AI systems, designing human-AI collaboration systems remains a critical problem in balancing efficiency with human agency.
+While agentic systems can autonomously execute actions on behalf of humans, their effectiveness hinges on aligning with user expectations and contextual needs.
+This challenge is particularly pronounced in domains requiring creativity and exploration.
+Taking image editing as an example, a task involving image parameter adjustment can yield numerous solutions that different creators come up with.
+Thus, agentic AI that automates the parameter adjustment can easily fall short of satisfying such subjective and aesthetic needs and lead to numerous iterations of prompt refinement.
+Instead, UIs with direct controls that bridge the communication between humans and agentic AI become essential.
+Thus, we argue that the user interface remains indispensable in agentic AI systems where meaningful human engagement is desired.
+Dataset injection to LLMs as a scalable solution for UI generation to support human-centered agentic AI: To allow interface interaction in agentic AI, LLM-generated UIs present a dynamic solution well-suited to the automation and adaptability required by LLM-based AI agents.
+However, using LLMs trained on generic data alone risks homogenizing design choices to generic solutions unless explicitly guided by target user-driven insights, as shown in this work.
+This highlights the need for UI generation frameworks to prioritize user needs.
+Among existing LLM post-training techniques for customizing LLMs to specific tasks and user preferences (Kumar et al., 2025; Tie et al., 2025), our framework contributes an approach of embedding a crowdsourced user dataset to guide LLMs to produce user-centric outputs through structured reasoning enhanced by this dataset.
+More interestingly, this framework can further benefit from agentic AI systems to deliver UI designs leveraging agent collaboration.
+Imagine that we can have multiple LLMs working simultaneously to generate UI designs, each equipped with user data covering different user profiles and design principles.
+A multi-agent system can be built to analyze these LLMs’ outputs, coordinate conflicts, and produce UI design suggestions with multiple perspectives considered.
+This agentic collaboration can enhance coverage and diversity in UI design, as well as support dynamic dialogue between agents to enable emergent design creativity.
 8. Limitations and Future Work
-The first limitation lies in the range of preferences covered during crowdsourcing.
-Our study focused on three key aspects, but there are other significant UI design principles, such as accessibility, cognitive load, and adaptability (Lidwell et al., 2010), that were not explicitly addressed here.
-However, our framework is highly extendable and can be enhanced by incorporating a broader set of preference dimensions with customizable aspects to account for user preferences from different perspectives.
-Next, we mainly considered software-based UI widget generation in this work.
-While these widgets suited our tasks, other potential options, like gesture-based or tangible widgets, were not included.
-This may limit our understanding of how different widgets, like those supporting spatial interactions in XR (Aponte et al., 2024) or physical interactions in tangible interfaces (Campos Zamora et al., 2024), could meet specific task needs or user preferences.
-Expanding the widget types in future research could offer comprehensive insights into user preference across a broader spectrum of tasks and interaction styles.
-Lastly, our system implementation is currently focused on UI generation within the image editing domain.
-In the future, we aim to collect data from multiple domains and various tasks to build a more extensive user preference library.
-This would allow us to evaluate how a cross-domain library can improve widget generation, both for in-domain tasks and for tasks outside the collected domains.
-Such a system would also introduce challenges, such as how to effectively manage large datasets and ensure the LLM fully leverages all the available knowledge.
+New interaction modalities beyond traditional GUIs: Although our work primarily focused on 2D UI widget generation, an emerging area is applying the framework to novel interaction modalities beyond 2D GUIs for preference alignment.
+For instance, as voice-controlled interfaces, AR/VR environments, and wearable device UIs become increasingly popular, it would be interesting to collect user preferences for those contexts.
+Future work can crowdsource user data regarding the voice command structure for a given function based on what most children find intuitive, gestures, or holographic controls that most senior people feel natural in AR, and gesture-based shortcuts or low-interruption interfaces for smartwatches suitable for most people in office settings.
+This direction can extend the crowdsourcing-informed interaction design as UIs diversify with technology advancements and user contexts.
+Whole-interface generation and layout: This work focuses on individual UI widget generation, and future research can expand to the complete interface layout or workflow generation by investigating how a crowdsourced dataset scales to decisions about arranging multiple widgets, navigation structures, or multi-step task flows.
+For example, given a complex task (e.g., “design a photo editing toolbar”), what interface layouts do users prefer, and how can an LLM assemble those using collected user insights?
+Crowdsourcing could be used to gather user-preferred layouts or sequences in addition to widget types.
+This can extend LLM-based UI generation to entire UIs that are human-centered at a macro level and also opens questions about representing and learning from more complex user data, e.g., sequences of interactions and grouping of controls, and ensuring coherence in generated designs.
+Integration with design guidelines and experts: Future work can combine the current crowdsourced general user data with established HCI design principles or input from professional UI designers.
+Research might explore hybrid systems where the LLM is guided by both the crowd-derived library and rule-based heuristics or where crowd preferences are augmented with expert critiques.
+This addresses whether certain nuanced design decisions (e.g., accessibility considerations or aesthetic consistency) might require more than just crowd preferences.
+By involving experts in the loop, e.g., crowdsourcing from designers or using expert-verified crowd data, the resulting UI generation could achieve a balance between general user appeal and professional best practices.
+Cross-cultural user data collection: The current work crowdsourced user preferences primarily from English-speaking participants.
+This can lead to overlooked preferences and cultural nuances that emerge in broader global contexts.
+For UI design, user preferences for color schemes, layout conventions, or even interaction metaphors may differ substantially across linguistic and cultural divides.
+Future data collection could employ multilingual crowdsourcing platforms to enhance cross-cultural inclusiveness and adapt tasks to reflect culturally relevant use cases.
+This would allow LLMs to account for cross-cultural insights to generate UIs that reflect the diversity of global user populations, thus promoting equitable and context-aware UI design on a global scale.
+User persona-aware UI design: Beyond user preferences, future research can incorporate broader user persona data, such as demographic information (e.g., age, gender), domain expertise (e.g., novice vs.
+expert users), device usage habits, cognitive styles, and other aspects, to represent essential aspects of target users.
+By leveraging these diverse persona attributes, the framework can be extended to process persona data and further generate UI designs that are not only preference-aligned but also contextually appropriate for different user types.
+Technical advancements of using LLMs for UI generation: This work implements a system prototype of the proposed framework and opens up opportunities for technical improvements.
+First, using LLMs for code generation introduces limitations, as minor bugs may occur and require manual fixes.
+Recent research has developed LLM-based agents to fix bugs in code automatically (Yang et al., 2024; Zhang et al., 2024; Kang et al., 2025).
+This offers a promising solution to robustly generate UIs in future iterations.
+Next, we use GPT-4o as the LLM for UI generation, and future research could explore more advanced LLMs for improved reasoning capabilities, open-source LLMs for lower cost and easier replicability, or fine-tuned LLMs for specific task domains.
+This would also lead to an in-depth comparison of different LLMs’ performance for UI generation.
 9. Conclusion
-In this paper, we introduced a UI widget preference library collected from a crowdsourcing study and a framework for LLM-driven widget generation enhanced by this library.
-Our evaluation of the framework demonstrated its ability to extend widget generation across diverse but related tasks, align generated widgets with user preferences captured through crowdsourcing, and use large libraries for consistent, preference-aligned widget reasoning while leveraging smaller libraries for cost-effective preference capture.
-We further discussed these results to explore the extendability of crowdsourcing tasks, preference aspects, and library sizes in supporting UI generation tailored to task-generic and task-specific use cases, diverse user preferences and needs, and adaptive workflows.
-We envision our framework to benefit both users and designers by simplifying the development of user-centered interfaces that enhance task efficiency and user satisfaction.
+In this paper, we introduced the CrowdGenUI framework for LLM-based widget generation enhanced by a crowdsourced user preference library.
+Through a user evaluation of the system prototype implemented for this framework, we demonstrated its generalizability to extend widget generation across diverse and related tasks, effectiveness to align generated widgets with user preferences, and capability to use large libraries for consistent preference-aligned widget reasoning and small libraries for cost-effective preference capture.
+We further discussed the user study results to explore how the framework can be extended to accommodate broader task domains, various user data, and scalable data sizes to support UI generation for emerging task contexts, diverse user needs, and adaptive, user-focused workflows.
+Based on the discussion, we suggested multiple future research directions regarding LLM-based UI generation and design.
+We envision our framework to benefit both end users and designers by simplifying the development of human-centered UIs that enhance task efficiency and user satisfaction.
 Appendix A LLM-Reasoned UI Widgets for User Evaluation Tasks
-Figures 13 to 18 show the LLM-reasoned UI widgets with crowdsourced libraries and without any library for the user evaluation tasks.
+Figures 12 to 17 show the LLM-reasoned UI widgets with crowdsourced libraries and without any library for the user evaluation tasks.
 Figure 12.
@@
 Appendix B Prompts for LLM UI Widget Reasoning and Code Generation
-B.1. Prompts for UI Widget Reasoning without Crowdsourced UI Preference Library
-Reason UI widget type for the user task. You should follow the steps below for the reasoning.
-First, please take the definitions below:
-- Predictability: allows users to obtain results with no surprises or doesnt require users to deduce how to perform the interaction.
-- Efficiency: allows users to perform tasks with a minimum amount of effort.
-- Explorability: allows users to explore multiple possibilities and perform functions with high flexibility.
-Second, reason the most proper UI widget for predictability, efficiency, and explorability.
-- The UI widget you reason must from Slider, Preset Buttons, Dropdown, Radio Buttons, Text Field, Color Wheel, Color Picker, Click on Image.
-Third, based on your reasoning, write down UI widgets for predictability, efficiency, and explorability in JSON format.
-Refer to the example below to provide your response.
-- Replace the placeholders marked by <> with your response. Do not include <> in your response.
-- You must keep all the existing information from the example and only replace placeholders.
-- The response must be in JSON format.
-"reasoning": {‘‘‘reasoning
-"task analysis": "<your reasoning>"
-"predictability_reasoning": {
-"<UI widget type>": "<your reasoning>"
-"efficiency_reasoning": {
-"<UI widget type>": "<your reasoning>"
-"explorability_reasoning": {
-"<UI widget type>": "<your reasoning>"
-"widget": {‘‘‘widget
-"<task_name>": {
-"predictability": "<UI widget type>",
-"efficiency": "<UI widget type>",
-"explorability": "<UI widget type>"
-B.2. Prompts for UI Widget Reasoning with Crowdsourced UI Preference Library
+B.1. Prompts for UI Widget Reasoning with Crowdsourced UI Preference Library
 Based on the crowdsourced UI widget preference library, reason UI widget type for the user task. You should follow the steps below for the reasoning.
@@
 "explorability": "<UI widget type>"
-B.3. UI Preference Library Appended to the UI Widget Reasoning Prompt
-Crowdsourced UI widget preference library task description:
-"image_adjust_lightness": "Adjust the lightness of the image.",
-"image_adjust_saturation": "Adjust the saturation of the image.",
-"image_adjust_hue": "Adjust the hue of the image.",
-"image_adjust_fall_color": "Change the image to fall colors.",
-"image_color_match": "Match the color of the image to a given target color.",
-"image_adjust_color_balance": "Adjust the color balance of the image by adjusting the RGB channels.",
-"image_place_watermark": "Place a watermark on the image to balance between visibility and subtlety.",
-"image_place_vignette": "Darken the background area using the vignette effect except for a certain image area."
-Crowdsourced UI widget preference library widget frequency:
-"image_adjust_lightness": {
-"Predictability": {
-"Slider": 1,
-"Preset Buttons": 4
-"Efficiency": {
-"Slider": 3,
-"Text Field": 1,
-"Radio Buttons": 1
-"Explorability": {
-"Slider": 5
-"image_adjust_saturation": {
-"Predictability": {
-"Slider": 2,
-"Preset Buttons": 3
-"Efficiency": {
-"Slider": 1,
-"Preset Buttons": 4
-"Explorability": {
-"Text Field": 1,
-"Slider": 4
-"image_adjust_hue": {
-"Predictability": {
-"Color Wheel": 4,
-"Preset Buttons": 1
-"Efficiency": {
-"Slider": 1,
-"Color Wheel": 1,
-"Preset Buttons": 2,
-"Radio Buttons": 1
-"Explorability": {
-"Color Wheel": 4,
-"Slider": 1
-"image_adjust_fall_color": {
-"Predictability": {
-"Color Wheel": 4,
-"Preset Buttons": 1
-"Efficiency": {
-"Preset Buttons": 2,
-"Slider": 1,
-"Radio Buttons": 1,
-"Color Wheel": 1
-"Explorability": {
-"Color Wheel": 4,
-"Slider": 1
-"image_color_match": {
-"Predictability": {
-"Text Field": 2,
-"Color Picker": 3
-"Efficiency": {
-"Text Field": 1,
-"Radio Buttons": 1,
-"Color Picker": 1,
-"Preset Buttons": 2
-"Explorability": {
-"Color Wheel": 4,
-"Color Picker": 1
-"image_adjust_color_balance": {
-"Predictability": {
-"Color Picker": 4,
-"Slider": 1
-"Efficiency": {
-"Color Picker": 3,
-"Text Field": 1,
-"Slider": 1
-"Explorability": {
-"Slider": 3,
-"Color Picker": 2
-"image_place_watermark": {
-"Predictability": {
-"Click on Image": 3,
-"Preset Buttons": 1,
-"Dropdown": 1
-"Efficiency": {
-"Click on Image": 2,
-"Preset Buttons": 3
-"Explorability": {
-"Preset Buttons": 1,
-"Click on Image": 2,
-"Slider": 2
-"image_place_vignette": {
-"Predictability": {
-"Preset Buttons": 2,
-"Click on Image": 2,
-"Slider": 1
-"Efficiency": {
-"Click on Image": 3,
-"Slider": 2
-"Explorability": {
-"Click on Image": 2,
-"Preset Buttons": 1,
-"Text Field": 1,
-"Slider": 1
-Crowdsourced UI widget preference library widget reasons:
-"image_adjust_lightness": {
-"Predictability": {
-"Slider": [
-"Its easy to understand that higher values will produce lighter images and viceversa"
-"Preset Buttons": [
-"The preset buttons background allows for users to figure out the lightness relatively quickly, and when clicking on the buttons, the button updates the lightness to how it looks on the buttons, allowing for users to need minimal deduction to understand what this button does.",
-"It is like to have an example of what you getting as a result",
-"The numbers show how each setting may look.",
-"the preset buttons provided an easy adjustment to the picture."
-"Efficiency": {
-"Slider": [
-"I can move it to the area I like the most",
-"Easier to adjust and more intuitive ",
-"the slider felt like the most efficient method to adjust the exposure. This method being something I am familiar with changing in Adobe programs."
-"Text Field": [
-"If a user were more experienced in photo editing, they can quickly adjust the lightness precisely to what they want it to be by simply typing in the value, making it quick and easy."
-"Radio Buttons": [
-"The values are displayed clearly and you just click one."
-"Explorability": {
-"Slider": [
-"Fast and contains options to the valid range so I dont have to figure that out.",
-"It allows a greater range with more fluid motion, it provides the result on the spot, and it feels natural. Can select more ranges from a wide variety to see different types of results.",
-"Allows for a more precise lighting value, and quicker",
-"Theres a bigger range of options",
-"I find there is a lot of sensitivity in the slider, which has greater impact."
-"image_adjust_saturation": {
-"Predictability": {
-"Slider": [
-"Saturation is usually set through a slider in image editing software",
-"using the slider was the easiest method to adjust the saturation. "
-"Preset Buttons": [
-"The preset buttons include a preview of how the saturation would look like.
-For example, -1.0 is fully gray while 1.0 is a bright rainbow.
-This allows for users who may not know what saturation means to understand what each button does, it changes how \"colorful\" the image is.",
-"It is easy to see the level of saturation that you want",
-"It shows what the colors are like and then I go from there."
-"Efficiency": {
-"Slider": [
-"Fast and easy to set it to a value I like."
-"Preset Buttons": [
-"The ability to view gradients and brightness on the preset buttons really gave me an advantage.
-I had an idea of the color spectrums brightness and what to expect, while the other options was almost like choosing in the dark and waiting for a surprise.",
-"Easier to use, and provides a glimpse into the saturation levels that radio buttons do not.",
-"You can just click the button and see the result assuming its a simple task like this one to just make colors pop",
-"This makes editing easier, as the values are predetermined."
-"Explorability": {
-"Text Field": [
-"I can enter specific values and see their results"
-"Slider": [
-"The slider allows users to slowly go through each increment and explore what effect they have on the image.
-They can go anywhere within the range to understand the effect of saturation on the image and precisely pinpoint which exact value of saturation they want with a slider.",
-"It is intuitive to increase the level of saturation just with the slider bar",
-"It discretely lists all possible colors and lets me just pick one while I scroll.",
-"the slider also contributed to the explorability."
-"image_adjust_hue": {
-"Predictability": {
-"Color Wheel": [
-"Colors dont feel like they should map to an order, and the color wheel is the only option that doesnt try to order the colors.",
-"For selecting the hue in this widget the color wheel was the most efficient choice in my opinion.
-It covers all the colors and gives me a clear area to select, the others I dont have much of guidance until I chose the options and see the results.",
-"Theres a lot more options and you know exactly what youre getting when you press an option.",
-"The color wheel provides more flexibility with color hues than the preset values."
-"Preset Buttons": [
-"The backgrounds of the preset buttons allow the user to already have a baseline as to what that hue value would then look like"
-"Efficiency": {
-"Slider": [
-"Im more familiar with using a slider to change the hue of an image"
-"Color Wheel": [
-"If the task were to match the image to the same hue as another image, the user could look at the reference image and quickly pinpoint it to which hue it matches closely to on the color wheel, without needing to explore additional options."
-"Preset Buttons": [
-"You obtain the colour that you choose",
-"the preset buttons provided the most efficient way to change the hue."
-"Radio Buttons": [
-"It gives a simple numeric representation which I can just click on."
-"Explorability": {
-"Color Wheel": [
-"I can select different colours and check the results",
-"There are more colour options ",
-"The color wheel shows how each color hue looks so I can see all combinations from the outside.",
-"the color wheel provided the most options to adjust the hue and explore other possibilities."
-"Slider": [
-"The slider allows for users to see the exact hue value as opposed to the color wheel, and see what even a small 0.01 increment in hue can do for an image.
-A slider for this task allows for more exploration as users scrub through the slider and see the different hue changes."
-"image_adjust_fall_color": {
-"Predictability": {
-"Color Wheel": [
-"Like before, colors dont have an order, so the wheel lets me know what I will get before clicking it.",
-"By being able to select a color on the wheel and it produce the widget result with that color I will say its the most appropriate option for predictability.
-I wouldnt have to deduce much knowing the style of orange or yellow I chose is what will be applied.",
-"Color wheel makes it easier to pick the colour you want, allowing for a more direct choice as you see the colours at once.",
-"I find it easer to navigate the wheel due to be able to seeing the colours beforehand."
-"Preset Buttons": [
-"Simple interface that has preset values and you know what to expecct"
-"Efficiency": {
-"Preset Buttons": [
-"I can quickly choose an option",
-"You can choose exactly the colour you want for the picture just with one click"
-"Slider": [
-"For this specific task, the slider is the most efficient option, because the other preset options do not create a autumnal atmosphere.
-The preset buttons only allow for red, where autumn is more orange or brown.
-The slider allowed for quickly scrubbing through each value to see what was most warm and orange, and then fine tuning the smaller 0.01 values for the exact atmosphere vibes."
-"Radio Buttons": [
-"You have few choices and can just click a simple choice."
-"Color Wheel": [
-"The color wheel was the most efficient way to perform the task. I was able to click on which color I thought I would like."
-"Explorability": {
-"Color Wheel": [
-"Its easier to test different possibilities and check the results",
-"The color wheel allows the user to explore different atmospheres for the image without having to figure out what 0-1.0 means in this context.
-The color wheel allows for the overall images atmosphere to be adjusted and explored in different seasonal contexts like pink for spring, and blue for winter.",
-"It lets you see the whole color spectrum and choose a particular hue that is right for the situation.",
-"The color wheel also offered more visual help to identify the colors and make the task more accessible."
-"Slider": [
-"As you can choose different values and see how it is affecting on the picture"
-"image_color_match": {
-"Predictability": {
-"Text Field": [
-"Since I know what my target it, just typing it gives me what I want.",
-"Can enter the hex number and see the desired result"
-"Color Picker": [
-"The color picker would accurately allow me to pinpoint the target color green and not have a doubt that it is correct.",
-"As you know what colour you are picking there are no surprises",
-"You can match the color visually before selecting it"
-"Efficiency": {
-"Text Field": [
-"Being able to directly type what I want is the faster than trying to use the slider or color picker or wheel."
-"Radio Buttons": [
-"It would be the easiest way to click the options while producing the results under minimal effort."
-"Color Picker": [
-"Allows you to directly choose the target colour, with minimal effort/attempts"
-"Preset Buttons": [
-"Simple interface with preset options. Little to no effort required",
-"This is much quicker to select, as the colors are preset"
-"Explorability": {
-"Color Wheel": [
-"Easy to try out options by just clicking.",
-"I feel like the wheel would have the most amount of options due to the color palette available to choose from.",
-"Allows for many options to be explored quicker than any of the other options",
-"This allows you to experiment more with the colors"
-"Color Picker": [
-"More explorability options and plenty of options to choose from."
-"image_adjust_color_balance": {
-"Predictability": {
-"Color Picker": [
-"Im familiar with using a colour picker to change the colours",
-"The color picker widget adjusts the overall images color harmony to the picked color quickly, and does not make the process confusing for the user to figure out how much red, green, and blue they need for the image to turn out a certain color.",
-"You need only to ckick the colour that you want to ",
-"The widget just lets you see a color and overlay it on the image."
-"Slider": [
-"I had no surprises using the slider to adjust the color balance."
-"Efficiency": {
-"Color Picker": [
-"It feels more intuitive",
-"It is easy to choose a colour, however there are less options than for the slider",
-"You can see what colors you want and just go with them without altering numbers."
-"Text Field": [
-"For more experienced users, the text field allows for them to efficiently type out the exact values they want to adjust without additional interaction.
-If the image needed just a little more blue, they could add 0.08 to it quickly, instead of fiddling with the slider."
-"Slider": [
-"The slider allowed me to perform the task with minimum effort. The slider felt more natural to use."
-"Explorability": {
-"Slider": [
-"Fastest way to just try different options.",
-"With multiple colors, the sliders makes it the most efficient and easy to use. It allows me to switch from colors easily and change their RGB values easily.",
-"Is the quickest out of all the options to explore many values other than the colour picker, however, the colour picker makes it harder to change individual colour channels."
-"Color Picker": [
-"Allows you click into colors and enter manually the RGB values.",
-"This allows users to explore more with the color settings"
-"image_place_watermark": {
-"Predictability": {
-"Click on Image": [
-"Clicking is both intuitive and I know what Ill get.
-It did take two clicks at least, because I thought the text would be centered on the click, but how it worked was quickly apparent so I could then adjust things as I wanted.",
-"As the user directly clicks on the image, they know exactly where the watermark will be placed.",
-"Text is there and you can just click on image and it will go exactly where you pressed."
-"Preset Buttons": [
-"The presets give an indication of what to expect and wouldnt confuse me in this task. I could choose the proper preset and its identifiable where the water mark would go."
-"Dropdown": [
-"It is easier to understand where the cursor will be positioned"
-"Efficiency": {
-"Click on Image": [
-"I can simply click where I want the watermark to be",
-"It is easier to choose the right position"
-"Preset Buttons": [
-"The typical locations a watermark would be placed in are listed in the preset buttons widget.
-In terms of efficiency, a user can quickly click on which location they want the watermark to appear and be done.
-This widget is also efficient in allowing the user to click through each option and explore where they want the watermark to be located quickly.",
-"They just give you a few options set out for you to pick from.",
-"The preset buttons are the most efficient way of adjusting the text."
-"Explorability": {
-"Preset Buttons": [
-"I can test different variations easily"
-"Click on Image": [
-"If a user wants more control over where their watermark is placed, the Click on image widget option is the best for this, because the user can click on exactly where they want the watermark to be, then fine tune it by clicking slightly, ultimately adjusting their watermark to their preferred location.",
-"You can totally visually discern where you want the watermarks and/or image to go to as it is continuous."
-"Slider": [
-"You can have more control by watching the numbers ",
-"The slider made the explorability more accessible to adjust the text."
-"image_place_vignette": {
-"Predictability": {
-"Preset Buttons": [
-"Its easier to understand where the circle will move to",
-"The preset buttons for this task allows for a user to quickly move the circle to common positions without any surprises.
-Users less familiar with photo editing can simply click on a button and know what it should be doing, and this widget does just that."
-"Click on Image": [
-"Very intuitive ",
-"I just click on a location and it goes where I want."
-"Slider": [
-"I feel that the slider came with no surprises. It was the easiest way to adjust the position of the circle."
-"Efficiency": {
-"Click on Image": [
-"The only option that really works well.",
-"Lets the user quickly determine where the circle will be placed",
-"You can point to exactly you need t be"
-"Slider": [
-"The easiest method would be the slider because it will produce the results with the least amount of effort. Its not the most powerful method but it is still efficient to a degree.",
-"It was the easiest to get the best outcome without taking a long time."
-"Explorability": {
-"Click on Image": [
-"Easy to just click in slightly different places as needed to get the circle where I want it.",
-"As it is the quickest method it also allows the user to check many positions more quickly than the other options."
-"Preset Buttons": [
-"The presets give me the most ranges with options to choose from that are already covering a wide variety. Its flexible and also easy to use."
-"Text Field": [
-"Allows you enter exact co-ordinates giving more flexibility."
-"Slider": [
-"This provides more flexibility with the positioning"
-B.4. Prompts for UI Widget Code Generation
+B.2. Prompts for UI Widget Code Generation
 Generate code for the UI widgets to perform the specified task. You should follow the steps below for the coding.
@@
 - If widget types and scores are the same, consider the reasons to make your selection.
-- All questions are required, including preference selection and an explanation of your reason.